### v1

# Causal machine learning for heterogeneous treatment effects in the presence of missing outcome data

Abstract
When estimating heterogeneous treatment effects, missing outcome data can complicate treatment effect estimation, causing certain subgroups of the population to be poorly represented. In this work, we discuss this commonly overlooked problem and consider the impact that missing at random (MAR) outcome data has on causal machine learning estimators for the conditional average treatment effect (CATE). We then propose two de-biased machine learning estimators for the CATE, the mDR-learner and mEP-learner, which address the issue of under-representation by integrating inverse probability of censoring weights into the DR-learner and EP-learner respectively. We show that under reasonable conditions, these estimators are oracle efficient, and illustrate their favorable performance through simulated data settings, comparing them to existing CATE estimators, including comparison to estimators which use common missing data techniques. Guidance on the implementation of these estimators is provided and we present an example of their application using the ACTG175 trial, exploring treatment effect heterogeneity when comparing Zidovudine mono-therapy against alternative antiretroviral therapies among HIV-1-infected individuals.
Key words: Causal machine learning; Influence functions; Heterogeneous treatment effects; Missing outcome data
1 Introduction
When evaluating the effect of an intervention, investigators are often interested in how the effect may vary within a target population. One approach used to explore treatment effect heterogeneity for a binary intervention is to estimate the conditional average treatment effect (CATE), defined as $\theta(x)=\operatorname{\mathbb{E}}[Y(1)|X=x]-\operatorname{\mathbb{E}}[Y(0)|X=x]$, where $Y(0)$ and $Y(1)$ are potential outcomes under the two levels of the treatment (Rubin, 2005) and $X$ represents the individual (pre-treatment) characteristics in which heterogeneity is of interest.
In recent years, many estimators for the CATE have been proposed, with the focus turning towards non-parametric estimators, using machine learning (ML) to estimate complex functions of high dimensional data (Künzel et al., 2019; Kennedy, 2023; Nie and Wager, 2021; van der Laan et al., 2024). Of these estimators, many require that the training data be fully-observed and no data be missing. In this paper, we explore how causal ML estimators can be constructed for the CATE when outcome data is missing at random (MAR), i.e., outcome missingness depends only on the treatment and observed baseline covariates.
MAR outcome data occurs frequently in practice, typically arising when individuals are lost to follow up. When it occurs, the observed data may no longer represent the target population, and subgroups that have high levels of drop-out can be under-represented. This presents a challenge for existing non-parametric CATE estimators, which do not address this under-representation and are prone to producing biased estimates of the CATE within these under-represented subgroups. To overcome this, some authors propose using CATE estimators in combination with established missing data techniques, such as imputing missing outcomes (Groenwold et al., 2014; Berrevoets et al., 2023), or re-weighting the population using inverse probability of censoring weights (IPCW) (Robins et al., 1994; Gonzalez Ginestet et al., 2021). However, when implementing these approaches using non-parametric (machine learning) techniques, the inherent slow convergence of ML algorithms (Hastie, 2009) can introduce errors into the IPCW or imputation predictions, which subsequently propagate through to the CATE estimates.
We explore this issue further, discussing the shortfalls of the existing CATE estimators when handling missing outcome data (Section 2). We then demonstrate how IPCWs can be robustly incorporated into CATE estimators, proposing two non-parametric CATE estimators, the mDR-learner and mEP-learner, which are extensions of the DR-learner (Kennedy, 2023) and EP-learner (van der Laan et al., 2024) respectively (Section 3). We demonstrate the empirical performance of our proposed estimators through a simulation study (Section 4) and illustrate their application to data from the ACTG175 trial (Hammer et al., 1996), where we explore treatment effect heterogeneity across baseline covariates, comparing Zidovudine mono-therapy against alternative antiretroviral therapies among HIV-1-infected subjects (Section 5). We then discuss the potential for further extensions in Section 6.
2 Background
2.1 Setting
We define a sample of $n$ independent and identically distributed (i.i.d.) observations drawn from $O=(Z,A,CY)$, where $A$ is a binary treatment, $A\in\{0,1\}$ and $Y$ is a continuous outcome. In this setting, we allow outcome data to be missing and define $C$ as an indicator of outcome missingness, $C\in\{0,1\}$, with $C=1$ indicating that an individual’s outcome is non-missing. We then define $Z$ to be a set of baseline covariates which contains all confounders between the treatment $A$ and the outcome $Y$, and also contains sufficient information for missing outcomes to be assumed MAR (i.e. after controlling for $Z$ and $A$, the missingness indicator $C$ and outcome $Y$ are independent).
The focus is on learning the CATE, $\theta(x)=\operatorname{\mathbb{E}}[Y(1)|X=x]-\operatorname{\mathbb{E}}[Y(0)|X=x]$, which is conditional on $X\subseteq Z$, where $X$ are the covariates for which heterogeneity is of interest.
2.2 Estimating the CATE using ML - No missing outcome data
Let us first consider how causal ML estimators are constructed when no outcome data is missing. Under the standard causal assumptions of (A1) consistency, (A2) no unmeasured confounding, $Y(a)\perp\!\!\!\perp A|Z$ for $a\in\{0,1\}$ and (A3) positivity of treatment exposure, $10<P(A=1|Z)<$ 1 (with probability 1) (Pearl et al., 2016), the CATE can be identified as:
$\theta(x)=\operatorname{\mathbb{E}}[\operatorname{\mathbb{E}}[Y|A=1,Z]-\operatorname{\mathbb{E}}[Y|A=0,Z]|X=x],$ (1)
or
$\theta(x)=\operatorname{\mathbb{E}}\left[\left.\frac{AY}{P[A=1|Z]}-\frac{(1-A)Y}{1-P[A=1|Z]}\right|X=x\right].$ (2)
ML can be used to directly estimate the conditional expectations or probabilities (nuisance functions) found in eq. (1) or eq. (2), with these estimators commonly referred to as plug-in estimators. One common example of a plug-in estimator is the T-learner (Künzel et al., 2019), which requires $X=Z$ and estimates the conditional expectations from eq. (1) in the subsets of individuals who are treated/untreated, $\mu^{1}(Z)=\operatorname{\mathbb{E}}[Y|A=1,Z]$ and $\mu^{0}(Z)=\operatorname{\mathbb{E}}[Y|A=0,Z]$, taking their difference to obtain CATE estimates. Alternatively, plug-in estimators based on eq. (2) are often referred to as inverse probability of treatment weight (IPTW) estimators, as they require the estimation of the propensity score, $\pi(Z)=P[A=1|Z]$ (Kennedy, 2023).
While such ML-based plug-in estimators are simple to understand and implement, they are also prone to bias. This arises as ML algorithms use regularization to avoid over-fitting to the training data, reducing the rate at which these estimators converge towards the true parameter. This introduces non-negligible bias when fit using finite samples, with the errors in these nuisance function estimates (e.g. $\hat{\mu}^{1}(Z),\hat{\mu}^{0}(Z)$ or $\hat{\pi}(Z)$) directly propagating through to the estimates of the CATE, known as “plug-in bias” (Morzywolek et al., 2023). Furthermore, plug-in estimators which estimate the conditional expectations of the outcome, such as the T-learner, do not ensure an optimal bias-variance trade-off is made for the CATE, as they optimize predictions for their outcome functions rather than for the CATE itself (Kennedy, 2023). See Appendix A for an illustrative example.
Estimators that target the CATE directly and that are less sensitive to nuisance model estimation avoid the issues noted above. Estimators of this form are typically constructed using the efficient influence function (EIF) of the estimand of interest, where the EIF represents how sensitive a measure of prediction error is to changes in the data generating distribution (Chernozhukov et al., 2017; Ichimura and Newey, 2022). The EIF offers a useful tool for constructing estimators, as it allows estimators to be designed which are approximately insensitive to small changes in its nuisance functions, meaning estimation errors in the outcome functions, $\mu^{1}(Z)$ and $\mu^{0}(Z)$, or propensity score, $\pi(Z)$, affect errors in the estimates of the target function only through their product. Unfortunately the EIF of the CATE is generally not well defined (i.e., has infinite variance) whenever it depends on continuous variables (Hines et al., 2022). Instead, estimators which achieve these properties can be constructed based on the EIF of a well-chosen loss function, in this case, a measure of counterfactual prediction error (Morzywolek et al., 2023). Two estimators constructed in this way are the DR-learner (Kennedy, 2023) and EP-learner (van der Laan et al., 2024).
The DR-learner is a model-agnostic method for estimating the CATE, meaning the user can choose any estimation strategy, including data adaptive methods, when estimating any functions within it. It is derived using a two-step procedure (Kennedy, 2023), with the first step calculating pseudo-outcomes, $Y_{DR}$, using the EIF of the mean square error (MSE) for the CATE:
$Y_{DR}=\frac{\left(A-\pi(Z)\right)}{\pi(Z)\left(1-\pi(Z)\right)}\left\{Y-\mu^{A}(Z)\right\}+\mu^{1}(Z)-\mu^{0}(Z),$ (3)
where $\hat{\mu}^{A}(Z)=A\cdot\hat{\mu}^{1}(Z)+(1-A)\cdot\hat{\mu}^{0}(Z)$. The second step then learns the CATE by regressing the pseudo-outcomes on the covariate set in which heterogeneity is of interest, $X$ (See Appendix B for the full algorithm).
The DR-learner is an example of a one-step IF based estimator and is derived by considering the components of the mean squared error (MSE) of the CATE that depend on $\theta(X)$,
$\psi_{CATE}=\operatorname{\mathbb{E}}\left[\theta(X)^{2}-2\theta(X)(\mu^{1}(Z)-\mu^{0}(Z))\right].$ (4)
It then uses the EIF of this risk function,
$\phi=\left(\theta^{2}(X)-\theta(X)\left(\mu^{1}(Z)-\mu^{0}(Z)\right)\right)+\frac{2\theta(X)(A-\pi(Z))}{\pi(Z)(1-\pi(Z))}\{Y-\mu^{A}(Z)\}-\psi_{CATE},$ (5)
and defines its pseudo-outcomes such that the sample average of the drift term (second term in eq. (5)) goes to 0. These pseudo-outcomes closely resemble the EIF from the average treatment effect (ATE) and remove the plug-in bias from the plug-in estimator, $\hat{\mu}^{1}(Z)-\hat{\mu}^{0}(Z)$, by adding the drift term, a weighted error, to these estimates. By defining the pseudo-outcomes in this way, the DR-learner ensures that the gradient of the MSE risk function, with respect to the CATE, is less sensitive to errors in its nuisance functions. It also allows the DR-learner to achieve oracle efficiency, meaning that when the outcome models and propensity score model converge to the truth at sufficiently fast rates, it performs asymptotically as if the nuisance functions were known (Kennedy, 2023). This allows data adaptive techniques to be used to estimate the nuisance functions, provided they converge sufficiently fast to the truth, without fear of errors propagating through to the CATE. However, we note that the convergence rate of the DR-learner will still depend on the convergence rate of its second-stage pseudo-outcome regression, which will vary itself depending on the complexity of the CATE.
Despite its desirable properties, the DR-learner’s pseudo-outcomes are sensitive to extreme propensity scores. This occurs as the DR-learner uses IPTWs when calculating its pseudo-outcomes (first term of (3)), meaning that when estimates of the propensity score are near 0/1, the pseudo-outcomes can grow infinitely large. To prevent this, an alternative approach known as infinite-dimensional targeting (iTMLE) can be used (Luedtke et al., 2017; Vansteelandt and Morzywołek, 2023; van der Laan et al., 2024). This technique has been developed for counterfactual outcome prediction (Vansteelandt and Morzywołek, 2023), and tends to have less sensitivity to extreme propensity scores as its targeted learning framework can moderate the impact of outlier propensity score estimates. In this paper we discuss a very similar approach known as the EP-learner (efficient plug-in learner) (van der Laan et al., 2024), which uses iTMLE in the context of causal contrasts, including CATE estimation.
The EP-learner is a model-agnostic estimator which also uses a two step procedure, first deriving pseudo-outcomes which are later regressed on $X$. However, the EP-learner uses an iTMLE procedure to generate its pseudo-outcomes. This iTMLE procedure plays a similar role to the one-step correction used in the DR-learner, and is motivated by the EIF of the MSE risk function for the CATE, eq. (5). Using this EIF, the iTMLE procedure aims to update the initial plug-in outcome estimates, $\hat{\mu}^{0}(Z)$ and $\hat{\mu}^{1}(Z)$ using targeted learning such that the sample average of the drift term (second term) in eq. (5), goes to 0. Yet, since the drift term in eq. (5) contains an infinite dimensional $\theta(X)$ (when any variables in $X$ are continuous), the classical targeted learning procedure which regresses outcomes against a scalar (known as a clever covariate) would fail to set the sample average of the drift term to 0. To resolve this, the iTMLE procedure defines a vector of univariate basis functions, referred to as a sieve, $\varphi(X)$, which it uses within the targeting step, approximating $\theta(X)$ and allowing the updates to the plug-in estimates to vary by $X$.
The targeting step then works by regressing the outcomes $Y$ on the sieve basis, $\varphi(X)$, in a weighted linear regression, with an offset $\mu^{A}(Z)$, and weight, $\hat{H}(A,Z)=\frac{A}{\hat{\pi}(Z)}+\frac{1-A}{\left(1-\hat{\pi}(Z)\right)}$, suggested by the EIF of the risk function, eq. (5). After fitting this model, efficient plug-in estimates, $\hat{\mu}^{0*}(Z)$ and $\hat{\mu}^{1*}(Z)$ are obtained by adding/subtracting the estimated linear predictor from this model to the plug-in outcome estimates (See Appendix C for the full algorithm). As the updated outcome predictions are defined such that the sample average of the drift term in eq. (5) converges to 0, the EP-learner achieves the same oracle efficiency properties as the DR-learner. Additionally, by using an iTMLE approach, extreme pseudo-outcomes are less common and the CATE estimates obtained by the EP-learner are expected to be more stable (van der Laan et al., 2024).
In addition, both the EP-learner and DR-learner require that the data which used within the second stage optimization is i.i.d. To achieve this, sample splitting is used when estimating the nuisance functions, preventing correlations from being introduced into the pseudo-outcomes. In this paper, we implement sample splitting by using a K-fold cross-validation process which works by splitting the data into K-folds, training each nuisance model on all folds except one, then obtaining predictions for the individuals in the remaining fold. This is repeated until nuisance function estimates are obtained for all individuals.
2.3 Estimating the CATE using ML - With missing outcome data
Now let us consider how missing outcome data impacts the existing estimators. To identify the CATE when outcome data is MAR, we require assumptions A1-A3 from Section 2.2 and two additional assumptions on the missingness mechanism: (A4) Outcomes are MAR, $Y\perp\!\!\!\perp C~{}|~{}A,Z$; (A5) Positivity of outcomes being non-missing, $0<P(C=1|A,Z)$, with probability 1.
Under assumptions A1-A5, the CATE can then be identified as eq. (6) and eq. (7).
$\theta(x)=\operatorname{\mathbb{E}}[\operatorname{\mathbb{E}}[Y|A=1,C=1,Z]-\operatorname{\mathbb{E}}[Y|A=0,C=1,Z]|X=x],$ (6)
or
$\theta(x)=\operatorname{\mathbb{E}}\left[\left.\frac{CAY}{P[C=1|A,Z]P[A=1|Z]}-\frac{C(1-A)Y}{P[C=1|A,Z](1-P[A=1|Z])}\right|X=x\right].$ (7)
The no unmeasured confounding assumption (A2) and MAR assumption (A4) need not be conditional on the same covariates, however, for simplicity, we define both assumptions to be conditional on the same set of covariates, $Z$.
Using equations (6) and (7) we see that when all of the covariates in $Z$ are present, the CATE can be estimated using the observed data. This means the existing CATE estimators, such as the T-learner, DR-learner and EP-learner can produce asymptotically unbiased estimates of the CATE by restricting their analyses to complete cases, and adjusting for $Z$ through their outcome models $\mu^{0}(Z)=\operatorname{\mathbb{E}}[Y|A=0,C=1,Z]$ and $\mu^{1}(Z)=\operatorname{\mathbb{E}}[Y|A=1,C=1,Z]$. We refer to this approach as an “available case analysis”, and note that while the outcome regressions are limited to complete cases, the propensity score models should still be estimated using the full dataset.
Available case analyses offer the simplest way of estimating the CATE in the presence of MAR outcome data, however, asymptotic unbiasedness does not mean these methods perform well when fit using finite samples, and available case analyses can be inefficient for a range of reasons. Firstly, as the outcome regressions are restricted to those with complete cases, information is ignored for individuals who have a missing outcome. Equally, when the outcome models are fit using ML, they will be prone to over-smoothing in the subsets of the population which experience high levels of outcome missingness. This can lead available case analyses to miss complex non-linear CATEs, over-smoothing outcome predictions due to the missing data, or can even lead to estimators identifying heterogeneity where non-exists, for instance when the missingness only occurs within one treatment arm. Specific examples of data generating processes (DGPs) where this can occur are presented in Section 4.
Because of the limitations of available case analyses, some authors choose to address missing outcome data by utilizing missing data techniques. One common missing data technique is to impute outcomes, replacing the missing outcomes with outcome predictions gained from an imputation model, $\hat{\operatorname{\mathbb{E}}}[Y|C=1,A,Z]$ (Groenwold et al., 2014; Berrevoets et al., 2023). These are easy to implement when using any of the existing non-parametric CATE estimators, as the estimators can be run using the imputed, complete dataset. However, as outcome imputations require the estimation of an additional nuisance model, plug-in bias can be introduced to the existing estimators, with errors in the outcome imputations propagating through to the estimates of the CATE (even for estimators which have oracle efficiency in the complete data setting). Alternatively, other authors suggest addressing missing outcome data by re-weighting observed individuals based on their probability of having a non-missing outcome, $G(A,Z)=P[C=1|A,Z]$ (Robins et al., 1994; Gonzalez Ginestet et al., 2021). These weights can be used to estimate the CATE using a similar approach to the one Kennedy (2023) uses to incorporate IPTWs, weighting observed individuals to create pseudo-outcomes, then regressing the pseudo-outcomes against $X$ to estimate the CATE. However, these estimators will also suffer from plug-in bias, as errors in the IPCWs will propagate through to the CATE estimates. For this reason, we explore how to incorporate IPCWs in both the DR-learner and EP learner such that the resulting learners are robust to estimation of the missingness model $G(A,Z)$.
3 DR-learner/EP-learner extensions
In this section, we extend the DR-learner (Kennedy, 2023) and EP-learner (van der Laan et al., 2024) to handle missing outcome data, leading to two new estimators, the mDR-learner (missing outcome DR-learner) and mEP-learner (missing outcome EP-learner) respectively.
3.1 mDR-learner
We begin by considering an extension of the DR-learner to the MAR outcome data setting. Recall that the DR-learner offers a one-step EIF based estimator of the CATE, constructing pseudo-outcomes using the EIF of the MSE for the CATE, which it regresses on the variables defining the CATE. By doing so, the DR-learner minimizes the MSE risk function and offers reduced sensitivity to errors in its nuisance functions, allowing data adaptive techniques to be used. This said, when outcomes are MAR, the DR-learner fit using the available case analysis does not account for the under-representation which occurs as a result of missing outcomes. This limitation occurs as the risk which the DR-learner uses to derive its pseudo-outcomes takes a new form when outcomes are MAR,
$\psi_{CATE}=\operatorname{\mathbb{E}}\left[\theta(X)^{2}-2\theta(X)(\operatorname{\mathbb{E}}[Y|C=1,A=1,Z]-\operatorname{\mathbb{E}}[Y|C=1,A=0,Z])\right],$ (8)
now involving the indicator for outcomes being non-missing. Consequently, the EIF of this risk function in this setting does not correspond to the EIF in the complete data setting. In Appendix D we derive the form of this EIF, showing it to be:
$\phi=\left(\theta^{2}(X)-2\theta(X)\left(\mu^{1}(Z)-\mu^{0}(Z)\right)\right)-\frac{2\theta(X)(A-\pi(Z))C}{\pi(Z)(1-\pi(Z))G(A,Z)}\{Y-\mu^{A}(Z)\}-\psi_{CATE}.$ (9)
Using this EIF, we can construct pseudo-outcomes for the mDR-learner which account for groups which are under-represented as a result of missing outcome data:
$Y_{mDR}=\frac{\left(A-\pi(Z)\right)C}{\pi(Z)\left(1-\pi(Z)\right)G(A,Z)}\left\{Y-\mu^{A}(Z)\right\}+\mu^{1}(Z)-\mu^{0}(Z).$ (10)
We note that IPCWs now appear in combination with IPTWs, allowing these pseudo-outcomes to account for the shift in covariate distribution caused by both missing outcome data and confounding. The mDR-learner then proceeds by regressing the estimates of these pseudo-outcomes $Y_{mDR}$ against covariates $X$ to obtain estimates of the CATE (See Algorithm 1).
As the mDR-learner defines its pseudo-outcomes using the EIF of the MSE for the CATE (under outcome missingness), it not only minimizes the MSE risk function, but also experiences less sensitivity to errors in its nuisance functions (including the IPCWs). We demonstrate this by exploring the excess risk of the mDR-learner, defined as the difference in MSE risk function, $L(.)$, when evaluated at $\hat{\theta}(X)$ and at $\theta^{*}(X)$, where $L(\theta^{*})=\inf_{\theta}L(\theta)$. Using the approach laid out by Foster and Syrgkanis (2023), we provide an upper bound for the excess risk (See Appendix E). This bound demonstrates how the gradient of the MSE risk function for the mDR-learner has reduced sensitivity to errors in the nuisance functions, which now include the missingness model, $G(A,Z)$. It also shows how the mDR-learner can obtain oracle efficiency under similar assumptions to those of the DR-learner.
For the mDR-learner to obtain oracle efficiency, it requires that the product of the errors from its outcome models and propensity score model, as well as the product of the errors from its outcome models and missingness model, converge at a faster rate than oracle estimates (i.e. estimates obtained from the second stage pseudo-outcome model under known nuisance functions). For instance, consider an example where the oracle CATE estimates converge at a $\sqrt{n}$ rate. In this example, if we wish to estimate the nuisance functions within the mDR-learner without impacting the overall convergence rate of our CATE estimates, then we require the estimators for each of our nuisance functions to converge at rates faster than $4]{n}$, hence allowing the product of errors to converge at a rate faster than $\sqrt{n}$. Equally, when the outcome function estimators converge at rates slower than this, oracle efficiency can still be obtained as long as the propensity score estimator and missingness function estimator converge at rates which are sufficiently fast for the above conditions to hold. Under these weakened convergence requirements, ML can then be used when estimating nuisance functions within the mDR-learner without errors propagating through to the CATE.
Algorithm 1 mDR-learner algorithm
1:Split the data randomly into $K$ (e.g., 10) equal sized folds of $n$ observations from $O=(Z,A,YC)$, denoted $D_{1},\ldots,D_{K}$.
2:For $j\in 1,...,K$ and using all folds $\{D_{i},i=1,\dots,K,i\neq j\}$ except $D_{j}$, train models for
[…]
3:For all individuals in $D_{j}$ ($j\in 1,...,K$), obtain predictions of $\hat{\pi},~{}\hat{G},~{}\hat{\mu}^{0}$ and $\hat{\mu}^{1}$, based on the models fitted in the remaining folds.
4:Construct the pseudo outcomes for each individual in the data using
$Y_{mDR}=\frac{\left(A-\hat{\pi}(Z)\right)C}{\hat{\pi}(Z)\left(1-\hat{\pi}(Z)\right)\hat{G}(A,Z)}\left\{Y-\hat{\mu}^{A}(Z)\right\}+\hat{\mu}^{1}(Z)-\hat{\mu}^{0}(Z)$
5:Regress the pseudo outcomes $Y_{mDR}$ on covariates $X$, and obtain predictions of $\theta(X)$:
$\hat{\theta}_{mDR}(X)=\hat{\operatorname{\mathbb{E}}}[Y_{mDR}|X]$
3.2 mEP-learner
We now demonstrate how the EP-learner, a targeted-learning based framework can be extended for the setting with MAR outcome data. As outlined in Section 2.2, this framework uses a two-step approach, first defining an iTMLE procedure to remove plug-in bias, then regressing the pseudo-outcomes obtained from this procedure against $X$ to obtain CATE estimates. However, when outcomes are MAR, we encounter the same problem discussed for the DR-learner, which is that the EIF used to derive its pseudo-outcomes now takes a new form, eq. (9). This EIF contains IPCWs within its drift term (second term), and hence, if we wish to set the sample average of the drift term to 0, removing plug-in bias, we must update the iTMLE process used within the EP-learner.
To update the iTMLE process, we re-define the weight using eq. (9): $\hat{H}(A,C,Z)=\frac{CA}{\hat{G}(A,Z)\hat{\pi}(Z)}+\frac{C(1-A)}{\hat{G}(A,Z)\left(1-\hat{\pi}(Z)\right)}$. By using this weight within the same iTMLE algorithm outlined previously, we ensure that the sample average of the second term in eq. (9) converges 0, which in turn means that the mEP-learner will be equivalent to the oracle efficient one-step mDR-learner when its nuisance function estimates converge sufficiently fast. The mEP-learner algorithm is provided in Algorithm 2, and for a general proof of this equivalence for EP risks, see the original EP-learner paper (van der Laan et al., 2024).
Algorithm 2 mEP-learner algorithm
1:Split the data randomly into $K$ (e.g., 10) equal sized folds of $n$ observations from $O=(Z,A,YC)$, denoted $D_{1},\ldots,D_{K}$.
2:For $j\in 1,...,K$ and using all folds $\{D_{i};i=1,\ldots,K,i\neq j\}$ except $D_{j}$, train models for
[…]
3:For all individuals in $D_{j}$ ($j\in 1,...,K$), obtain predictions of $\hat{\pi},~{}\hat{G},~{}\hat{\mu}^{0}$ and $\hat{\mu}^{1}$, based on the models fitted in the remaining folds.
4:For all individuals in the data, update the outcome predictions
  1. a)
    Calculate the clever covariate, $\hat{H}(A,C,Z)=\frac{CA}{\hat{G}(A,Z)\hat{\pi}(Z)}+\frac{C(1-A)}{\hat{G}(A,Z)\left(1-\hat{\pi}(Z)\right)}$.
  2. b)
    Choose a sieve basis, $\hat{\varphi}(X)$
  3. c)
    Run a linear regression of outcomes $Y$ on feature vector $\hat{\varphi}(X)$ with offset $\hat{\mu}^{A}(Z)$ and weight $\hat{H}(A,C,Z)$ in the complete cases.
  4. d)
    Estimate the coefficients $\hat{\epsilon}$ from c) and use these to update $\hat{\mu}^{0}(Z)$ and $\hat{\mu}^{1}(Z)$
    $\hat{\mu}^{1*}(Z)=\hat{\mu}^{1}(Z)+\hat{\epsilon}\cdot\hat{\varphi}(X),~{}~{}\hat{\mu}^{0*}(Z)=\hat{\mu}^{0}(Z)-\hat{\epsilon}\cdot\hat{\varphi}(X)$
5:Construct the pseudo outcome for all individuals in the data
$Y_{mEP}=\hat{\mu}^{1*}(Z)-\hat{\mu}^{0*}(Z)$ (19)
6:Regress the pseudo outcomes $Y_{mEP}$ on covariates $X$, and obtain predictions of $\theta(X)$:
$\hat{\theta}_{mEP}(X)=\hat{\operatorname{\mathbb{E}}}[Y_{mEP}|X]$
3.3 Implementation
As both the mDR-learner and mEP-learner serve as general frameworks for estimating the CATE, their implementation requires the user to make several key decisions. In this section, we break these decisions down into two groups; decisions required for obtaining CATE estimates, and decisions required for assessing CATE performance.
3.3.1 Decisions required for obtaining CATE estimates
To obtain CATE estimates from either the mDR-learner or mEP-learner, the user must first decide which estimation technique they will use to estimate the nuisance functions and second stage model. As both learners are model-agnostic, this choice is flexible, and data adaptive techniques can be used. However if oracle performance is to be achieved, the estimators from the nuisance models must converge sufficiently fast to the truth. For this reason, we illustrate their implementation using the Super Learner (Van der Laan et al., 2007), an ensemble learner that allows a range of data-adaptive algorithms to be implemented and which performs asymptotically as well as its best candidate learner.
After algorithm choice, users are also required to decide which type of sample splitting they will implement within the mDR or mEP learners, ensuring the data used in the pseudo-outcome regression is i.i.d. Various options exist for achieving this, including K-fold cross-validation (see Section 2.2) or independent sample splitting, which splits the data in two before training the nuisance models/second stage model on each half separately (Kennedy, 2023). However, estimators that allocate fully independent data for each nuisance/target model are typically less efficient, with the smaller sample sizes introducing finite sample bias into the estimation process. For this reason, we demonstrate the application of the mDR-learner and mEP-learner by using a 10-fold cross-validation process (see Algorithm 1), however, we note that the appropriate number of folds will depend on the complexity and smoothness of the underlying nuisance/target functions. If users wish to explore variations of cross-fitting approaches, sensitivity analyses could be run using alternative numbers of splits.
We also highlight that when cross-validation is used, the reduction in training sample size for each model can introduce positivity violations, leaving certain subgroups poorly represented within some folds. To overcome this, the mDR-learner and mEP-learner can be run a large number of times ($J$ times), using a different seed for the cross-fitting split. This results in a vector of CATE estimates for each individual, $\hat{\theta}_{s}(X)$, $s=1,..,J$. Final CATE estimates are then obtained by taking the median across these estimates, $\widehat{\theta}(X)=median\{\widehat{\theta}_{s}(X)\}$ (Chernozhukov et al., 2018; Jacob, 2020).
Finally, when implementing the mEP-learner, the sieve basis used within the iTMLE process must be specified. The existing iTMLE implementations (Vansteelandt and Morzywołek, 2023; van der Laan et al., 2024) use a univariate trigonometric cosine polynomial basis, as it offers strong approximation guarantees under smoothness assumptions (Zhang and Simon, 2023). We also implement this sieve basis, following the guidance of Zhang and Simon (2023) to define the dimension of the sieve and its default interaction order. However, a cross-validation option for choosing the tuning parameters of this sieve was presented by van der Laan et al. (2024). In addition, we note that when implementing the iTMLE procedure, penalization methods, rather than a linear regression, can be used in the updating step. This was demonstrated by (Vansteelandt and Morzywołek, 2023), and can allow for better model performance when the sieve grows very large.
3.3.2 Decisions required for assessing CATE performance
The decisions outlined above enable users to obtain CATE estimates from the mDR-learner and mEP-learner. In practice, however, users may also want to assess the accuracy of these estimates by obtaining measures of uncertainty (e.g., confidence intervals) or calculating evaluation metrics such as the mean squared error (MSE).
Calculating measurements of uncertainty for CATE estimates generated using non-parametric estimators is challenging, as the theoretical convergence guarantees required for confidence intervals to be derived are often not met when using non-parametric estimation techniques. However, recent work by Takatsu and Westling (2022); Bonvini et al. (2023) and Ritzwoller and Syrgkanis (2024) offer some solutions. Bonvini et al. (2023) provides options for obtaining confidence intervals when kernel based approaches are used to estimate a univariate, additive or has partial dependence CATE, while Ritzwoller and Syrgkanis (2024) offers a half-sample bootstrapping approach which can provide confidence intervals when kernel based approaches are used to estimate the pseudo-outcome regression. As the work we present in this paper is focused on outlining general frameworks for estimating the CATE in the presence of MAR outcome data, we leave the application of these novel approaches to the user.
We also note how evaluating the performance of CATE estimators in real (as opposed to simulated) data can be problematic, as only one of the two potential outcomes, $Y(0)$ and $Y(1)$, is ever observed for each individual. Consequently, calculating standard metrics such as the MSE in practice is not possible, and instead, estimators of these metrics must be constructed. Simple estimators of the MSE, using either potential outcome predictions on their own, or in combination with IPTWs have been proposed (Schuler et al., 2018), however, these estimators are not robust to errors in the outcome or propensity score models and are prone to plug-in bias. Equally, EIF based estimators which are typically used to address plug-in bias such as this, encounter their own issues, as by definition, the MSE metric and its EIF will be 0 when CATE estimates are unbiased. Alternatively, model performance can be assessed by reviewing the stability of model estimates across different sample splitting seeds, enabling users to understand the variability in individual CATE estimates obtained from estimators such as the mDR-learner or mEP-learner.
4 Simulation study
4.1 Set up
We study the finite sample performance of the mDR-learner and mEP-learner across three data generating processes (DGPs), where each DGP corresponds to a setting in which missing outcome data can complicate the estimation of the CATE.
In each setting, we generate 6 uniformly distributed covariates $Z$ (independently), a binary treatment $A$ and a continuous outcome $Y$. In the first two DGPs, we define a simple unexposed outcome function $\mu^{0}(Z)$, a complex CATE $\theta(X)$ and define outcome missingness such that it occurs with high probability in only the treatment arm (DGP 1), or in both arms (DGP 2). This missingness mechanism makes the complex CATE more challenging to learn. In the third DGP, we define a complex unexposed outcome function $\mu^{0}(Z)$, a simple CATE $\theta(X)$ and allow outcome missingness to occur evenly across treatment arms, now making the simple CATE difficult to learn. More details can be found in Appendix G.
We vary the training data sample size from 400 to 3200, and use 500 replicates for each scenario. We compare the mDR-learner and mEP-learner to three alternative CATE estimators; the DR-learner, EP-learner and the T-learner, each implemented using (1) available cases (Section 2.3) and (2) imputed outcomes, obtained from imputation models. All estimators were implemented using 10 fold cross-validation and all models (including imputation models) were fitted using the Super Learner (Van der Laan et al., 2007), which included linear regression, LASSO, random forest and neural nets, amongst other algorithms in its library. Further details, including tuning parameters are reported in Appendix H.
To assess the performance of each estimator, we generated one test dataset with sample size $n=10,000$ per DGP. For each estimator applied to each DGP we obtained the CATE estimates for each individual in this test dataset. Performance was measured by calculating the root mean square median error (RMSME) of each learner (Appendix I). We use this metric as mean root mean square error (RMSE) estimates were found to be skewed when using 500 replications. For comparisons made using the mean RMSE, see Appendix J.
4.2 Findings
Figure 1 reports the RMSME for each candidate learner. The left column compares the mDR-learner and mEP-learner with the T-learner, DR-learner and EP-learner using available cases. In this comparison we see that the mDR-learner has better performance than the DR-learner across all three DGPs. The same trend can also be seen for the mEP-learner when compared with the EP-learner. The right-hand column shows the performance of the T-learner, DR-learner and EP-learner using imputed outcomes. Here, the mDR-learner and mEP-learner outperformed their corresponding learner with imputed outcomes when the CATE was complex (DGP 1 and 2), demonstrating that the errors introduced by imputing the outcomes propagated through the CATE estimators. However, when the CATE was simple, the DR-learner and EP-learner combined with outcome imputations performed well.
Compared to the available case/imputed outcomes T-learner, the mDR-learner and mEP-learner were less sensitive to specific nuisance function specification, while the T-learner’s performance depends greatly on the complexity of the underlying outcome functions, and worsens substantially when the outcome functions are complex (DGP 3).
Finally, the CATE estimates obtained across simulations for the mEP-learner, EP-learner and T-learner were more stable than those obtained by the mDR-learner and DR-learner. In both the mDR-learner and DR-learner, extreme CATE estimates were produced in some simulated settings, as positivity violations within the folds of the training data occurred. This can be seen more clearly when performance is measured using mean RMSE (Appendix J).
Figure 1: Root mean square median error (RMSME) for mDR-learner, mEP-learner, DR-learner, EP-learner and T-learner in three DGPs plotted by training sample size. Plots in the left column compare the mDR-learner and mEP-learner to the available case versions of the DR-learner, EP-learner and T-learner in DGP 1, 2 and 3 respectively. Plots in the right column compare the mDR-learner and mEP-learner to the DR-learner, EP-learner and T-learner when used in combination with an outcome imputation model in DGP 1, 2 and 3 respectively.
5 ACTG175 trial analysis
5.1 Background
We illustrate the use of the mDR-learner and mEP-learner by applying them to the ACTG175 trial (Hammer et al., 1996), which studied the effect of Zidovudine mono-therapy vs alternative antiretroviral therapies among HIV-1-infected subjects. The ACTG175 trial is a randomised, double blind, placebo controlled trial which recruited 2139 patients from 43 AIDS Clinical Trial Units and 9 National Hemophilia Foundation sites in the United States and Puerto Rico, from 1991 to 1994. Participants were randomised to one of four regimens: Zidovudine mono-therapy; Zidovudine and Didanosine; Zidovudine and Zalcitabine or Didanosine mono-therapy, with the randomisation stratified according to the length of prior antiretroviral therapy.
Baseline covariates were recorded including information on demographics, medical history and disease progression. Lab measurements of CD4 and CD8 T-cell counts (cells/mm3) were taken at baseline as well as at $20\pm 5$ weeks and $96\pm 5$ weeks, with the measurement at $96\pm 5$ weeks defined as the end point of interest. Of the 2139 patients randomised, 797 (37%) were lost to follow-up, either dropping out or dying during the follow up period, however, information on reason for being lost to follow-up is not present in the available data.
We focus on an intention-to-treat comparison between Zidovudine mono-therapy ($n=532$) and all alternative therapy lines ($n=1607$), as done by Wang et al. (2021). We estimate two CATEs, one conditional on all baseline covariates, and one conditional only on age (Fan et al., 2017).
5.2 Methods
CATEs are estimated using the mDR-learner and mEP-learner, and compared to estimates from the DR-learner, EP-learner and T-learner, all fit using available case analyses (Section 2.3). Propensity score, missingness and outcome models, all include the following 12 baseline covariates: age (years), gender (male/female), race (white/non-white), weight (kg), hemophilia (yes/no), homosexual activity (yes/no), Karnofsky score (on a scale of 0-100), symptomatic status (symptomatic/asymptomatic), history of intravenous drug use (yes/no), antiretroviral history (experienced/naive), CD4 T-cell count (cells/mm3) at baseline and CD8 T-cell count (cells/mm3) at baseline. All nuisance models are implemented with a Super Learner and details on the libraries/tuning parameters used within the Super learner are outlined in Appendix H. All methods are implemented using 10 fold cross-validation and we report the median CATE estimate over 50 different sample splitting random seeds.
5.3 Findings
Forty percent of Zidovudine mono-therapy participants were lost to follow up by week 96+/-5, while the alternative arm had 36% lost to follow up. The covariate distribution of the observed patients in each arm was not substantially different from the recruited trial participants, only differing slightly in CD4/CD8 T-cell counts at baseline. However we note that the observed mono-therapy arm appeared to be healthier than the recruited population, suggesting that individuals with lower CD4/CD8 cell counts in this arm were more likely to drop-out. In addition, the sample size in each arm among individuals $<20$ years old and $>60$ years old were both very small, with each group in each arm representing less than $2\%$ of their observed populations.
Figure 2 shows the median CATE estimates (over 50 different random seeds for sample splitting), conditional on age only. In this setting, we obtained similar results between the mDR-learner and the available case DR-learner, both suggesting a constant treatment effect between the ages of $20-60$, and an uptick of efficacy for alternative antiretroviral therapies amongst those under 20 years old and over 60 years old. Meanwhile, the mEP-learner and available-case EP-learner also performed comparably, instead suggesting a smoother treatment effect curves over all ages. We note that the difference in the results for those in the lower/higher age range may arise from the poor representation at these ages, highlighting the DR-learner and mDR-learner’s greater sensitivity to positivity violations compared with the EP-learner and mEP-learner. This can be seen by looking at learner estimates across a range of random splits used to fit the data (Figure 3).
We also consider the CATE estimates conditional on all baseline covariates, with Table 1 reporting the mean CATE estimates for individuals in five age groups. Once again, the estimates were comparable within learner type, however, we see the T-learner estimated smaller treatment effects across all ages, potentially demonstrating that the T-learner fails to account for the healthier observed population in the mono-therapy control group.
Figure 2: CATE estimates from mDR-learner, DR-learner and T-learner plotted by age (years).
Figure 3: CATE estimates from mDR-learner, DR-learner and T-learner plotted by age (years) for 4 different seeds.
Table 1: Mean CATE estimates by age groups when allowing the CATE to be conditional on all baseline covariates.
Age (Categorical) mDR-learner mEP-learner DR-learner EP-learner T-learner
$<$ 20 89.730 53.917 80.846 53.639 59.940
20-29 62.601 59.185 61.300 59.173 57.980
30-39 63.109 63.255 62.319 63.115 62.473
40-49 62.588 66.004 62.322 65.731 63.292
$\geq$ 50 64.174 70.672 64.776 70.360 62.033
6 Discussion
In this paper, we discussed the commonly overlooked problem of estimating the CATE when outcome data is MAR. Our work proposes two robust oracle efficient estimators, the mDR-learner and mEP-learner, which extend the DR-learner and EP-learner respectively, incorporating IPCWs and allowing the learners to address the population imbalances introduced by missing outcome data while maintaining favorable convergence properties. We demonstrated their robust performance through a simulation study and an applied data example, and provide guidance on the different decisions required in the estimation process.
Our proposed approaches have the potential to help improve CATE estimation in real-world data settings where outcome data are missing. However, there remains considerable scope for further developments of methods for the estimation and evaluation of CATEs. For example, further extensions of these techniques to handle more complex data, e.g., post-baseline covariate information or missing covariate data would allow for these techniques to see wider use. We outline an example of one of these extensions in Appendix K, where we discuss how the mDR-learner could be extended to handle post-baseline covariates which impact outcome missingness over time. Additionally, although we present potential options for measuring uncertainty or evaluating non-parametric CATEs in Section 3.3, further work on understanding when these techniques work well would greatly improve the utility of the estimators presented in this paper.
Acknowledgements
This work was supported by the Medical Research Council [grant number MR/N013638/1]. KDO was funded by a Royal Society-Welcome Trust Sir Henry Dale fellowship, grant number 218554/Z/19/Z. RHK was funded by UK Research and Innovation (Future Leaders Fellowship MR/X015017/1).
Appendix A Illustration of T-learner bias
We illustrate, through simulated data, how the T-learner, which is constructed by optimizing predictions for conditional mean outcome functions in the treated and untreated arms, does not deliver the correct bias-variance trade-off.
Consider three covariates $\{Z_{1},Z_{2},Z_{3}\}$, generated uniformly between -1 and 1. Then let $A$ be a binary treatment from a Bernoulli distribution with probability of $\frac{1+sin(Z_{1})}{2}$, and $Y$ be a binary outcome $Y$ with probability:
$\frac{1}{1+e^{-\left(\frac{4Z_{1}^{2}}{2}\cdot\mathbbm{1}(Z_{1})-Z_{1}\cdot(1-\mathbbm{1}(Z_{1}))+Z_{1}\cdot(1-\mathbbm{1}(Z_{1}))+\frac{Z_{2}}{2}+\frac{Z_{3}}{2}\right)}}$
where
$\mathbbm{1}(Z_{1})=\begin{cases}0&x<-0.5\\ 1&-0.5\leq x\leq 0.5\\ 0&0.5<x.\end{cases}$
Note that in this example no treatment effect exists, i.e. the difference between the two functions $\mu^{1}$ and $\mu^{0}$ is a simple function (a constant, equal to zero), but the individual functions are complex and therefore difficult to estimate.
Figure 4 presents 2000 data points from this example, with the true $\mu^{1}$ and $\mu^{0}$, i.e. the logit of the potential outcome probabilities plotted against their $Z_{1}$ values, for untreated and treated individuals.
Figure 4: Simulated data example plotting the logit of the outcome probabilities in each treatment arm across $Z_{1}$ values for 2000 data points. Dashed line presents the estimates made by the T-learner.
In this setting, the untreated individuals with $Z_{1}$ values above 0.5 are under-represented, while treated individuals with $Z_{1}$ values less than -0.5 are also under-represented. As a result, the outcome models for the treated/untreated are more difficult to fit in those regions, and the machine learning models needed to fit a T-learner will over-smooth in the respective areas of $Z_{1}$ which are under-represented. Figure 4 demonstrates this by plotting the estimated $\hat{\mu^{a}}(Z)$ obtained using tuned random forests, across 500 simulations, which are later used to construct a T-learner.
As a consequence of the uneven smoothing across the treated/untreated groups, the resulting plug-in CATE estimates are volatile and heavily prone to bias (Figure 5). In contrast, as the DR-learner and EP-learner which target the CATE directly, express less volatility.
Figure 5: Mean of CATE estimates across 500 simulations, plotted by $Z_{1}$ values for the T-learner, DR-learner and EP-learner.
Appendix B DR-learner algorithm
Algorithm 3 DR-learner algorithm
1:Split the data randomly into $K$ (e.g., 10) equal sized folds of $n$ observations from $O=(Z,A,Y)$. Denote $D_{1},\ldots,D_{K}$.
2:For $j\in 1,...,K$ and using all folds $\{D_{i},i=1,\dots,K,i\neq j\}$ except $D_{j}$, train models for
[…]
3:For all individuals in $D_{j}$ ($j\in 1,...,K$), obtain predictions of $\hat{\pi},~{}\hat{\mu}^{0}$ and $\hat{\mu}^{1}$, based on the models fitted in the remaining folds.
4:Construct the pseudo outcomes for each individual in the training data using
$Y_{DR}=\frac{A-\hat{\pi}(Z)}{\hat{\pi}(Z)\left(1-\hat{\pi}(Z)\right)}\left\{Y-\hat{\mu}^{A}(Z)\right\}+\hat{\mu}^{1}(Z)-\hat{\mu}^{0}(Z)$
5:Regress the pseudo outcomes $Y_{DR}$ on covariates $X$, and obtain predictions of $\theta(X)$:
$\hat{\theta}_{DR}(X)=\hat{\operatorname{\mathbb{E}}}[Y_{DR}|X]$
Appendix C EP-learner algorithm
Algorithm 4 EP-learner algorithm
1:Split the data randomly into $K$ (e.g., 10) equal sized folds of $n$ observations from $O=(Z,A,Y)$. Denote $D_{1},\ldots,D_{K}$.
2:For $j\in 1,...,K$ and using all folds $\{D_{i},i=1,\dots,K,i\neq j\}$ except $D_{j}$, train models for
[…]
3:For all individuals in $D_{j}$ ($j\in 1,...,K$), obtain predictions of $\hat{\pi},~{}\hat{G},~{}\hat{\mu}^{0}$ and $\hat{\mu}^{1}$, based on the models fitted in the remaining folds.
4:Using all individuals in the data, update outcome predictions
  1. a)
    Construct the clever covariate, $\hat{H}(A,Z)=\frac{A}{\hat{\pi}(Z)}+\frac{1-A}{\left(1-\hat{\pi}(Z)\right)}$.
  2. b)
    Choose a sieve basis, $\hat{\varphi}(X)$
  3. c)
    Run a linear regression of outcomes $Y$ on feature vector $\hat{\varphi}(X)$ with offset $\hat{\mu}^{A}(Z)$ and weight $\hat{H}(A,Z)$ in the complete cases.
  4. d)
    Estimate the coefficients $\hat{\epsilon}$ from c) and use these to update $\hat{\mu}^{0}(Z)$ and $\hat{\mu}^{1}(Z)$
    $\hat{\mu}^{1*}(Z)=\hat{\mu}^{1}(Z)+\hat{\epsilon}\cdot\hat{\varphi}(X),~{}~{}\hat{\mu}^{0*}(Z)=\hat{\mu}^{0}(Z)-\hat{\epsilon}\cdot\hat{\varphi}(X)$
5:Construct the pseudo outcome for all individuals in the training data
$Y_{EP}=\hat{\mu}^{1*}(Z)-\hat{\mu}^{0*}(Z)$ (26)
6:Regress the pseudo outcomes $Y_{EP}$ on covariates $X$, and obtain predictions of $\theta(X)$:
$\hat{\theta}_{EP}(X)=\hat{\operatorname{\mathbb{E}}}[Y_{EP}|X]$
Appendix D Derivation of the EIF used within the mDR-learner and mEP-learner
Let us define $\mathcal{P}$ as the true observed data distribution and let our estimand of interest be
$\psi(\mathcal{P})=\operatorname{\mathbb{E}}_{\mathcal{P}}[\theta(X)^{2}-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}(Y(1)|X)-\operatorname{\mathbb{E}}_{\mathcal{P}}(Y(0)|X))].$
Assuming A1-A5 hold, we can write this risk function as:
$\psi(\mathcal{P})=\operatorname{\mathbb{E}}_{\mathcal{P}}\left[\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z]-\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=0,C=1,Z])\right]$
Expanding this, we can write our estimand as
[…]
To derive the efficient influence function for (27), we perturb $\mathcal{P}$ in the direction of a point mass at $(\tilde{z},\tilde{a},\tilde{c})$ to get:
[…]
where,
$f_{t}(z)=t\mathbbm{1}_{\tilde{z}}(z)+(1-t)f(z)$
We can then derive the EIF by calculating the Gateaux derivative:
$\frac{d\psi(\mathcal{P}_{t})}{dt}\Bigr{|}_{t=0}$ $=\int\frac{d}{dt}\left\{f_{t}(z)\left(\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=1,C=1,Z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=0,C=1,Z])\right)\right\}\Bigr{|}_{t=0}dz$
=$=\int\left\{(\mathbbm{1}_{\tilde{z}}(z)-f(z))\left(\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=0,C=1,Z])\right)+\right.$
$~{}~{}~{}~{}~{}~{}~{}\left.f(z)\frac{d}{dt}\left\{\left(\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=1,C=1,Z]-\right.\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=0,C=1,Z])\right)\right\}\right\}|_{t=0}dz$
=$=\int\left\{(\mathbbm{1}_{\tilde{z}}(z)-f(z))\left(\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=0,C=1,Z])\right)+\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}f(z)\left(-2\theta(X)\frac{d}{dt}(\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=1,C=1,Z]-\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=0,C=1,Z])|_{t=0}\right)\}dz$
=$=\int\left\{(\mathbbm{1}_{\tilde{z}}(z)-f(z))\left(\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=0,C=1,Z])\right)+\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}f(z)\left(-2\theta(X)\frac{\mathbbm{1}_{\tilde{z}}(z)(A-\pi(Z))C}{f(z)(1-\pi(Z))\pi(Z)G(A,Z)}\{y-\mu_{A}\}\right)\}dz$
$=\theta(X)^{2}-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z]-\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z])-$
$~{}~{}~{}~{}2\theta(X)\frac{(A-\pi(Z))C}{(1-\pi(Z))\pi(Z)G(A,Z)}\{y-\mu_{A}\}-\psi(\mathcal{P})$
where $\pi(Z)=f(A=1|Z)=P(A=1|Z=z)$, $G(A,Z)=f(C=1|A,Z)=P(C=1|A,Z)$ and $\mu^{a}(Z)=\operatorname{\mathbb{E}}_{\mathcal{P}}(Y|A=a,C=1,Z)$
Appendix E mDR-learner excess risk bound - Overview
To demonstrate that the estimation error of the nuisance functions only has a second-order impact and therefore the mDR-learner is approximately insensitive to this, we provide an upper bound for the excess risk of the mDR-learner. We do so by following the structure and Theorems laid out by Foster and Syrgkanis [2023].
We define the excess risk as the difference
$L(\hat{\theta})-L(\theta^{*})$ (28)
in MSE risk $L(.)$ when evaluated at $\hat{\theta}$ and at $\theta^{*}$, where $\theta^{*}$ is defined as the minimiser of risk function when the infimum of the risk function is obtained, $L(\theta^{*})=\inf_{\theta}L(\theta)$.
Let us also define $\Theta$ be the class of target parameter values, and $\hat{\Theta}$ to be the class of target parameter estimates, containing all possible $\hat{\theta}$. As the population risk function depends not only on $\theta\in\Theta$, but on a set of nuisance parameters $g=(\pi,G,\mu^{0},\mu^{1})$, we introduce to notation $L(\theta,g)$. Additionally, as the objective of the mDR-learner is to minimise the MSE risk function under known nuisance functions $g_{0}$, we write the excess risk that we wish to bound as:
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0}).$ (29)
When the nuisance functions are known, an upper bound for this excess risk can be written in terms of the convergence rate of the CATE estimates obtained from mDR-learner’s second stage pseudo-outcome regression:
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})\leq\text{Rate}_{\theta}(g_{0}).$
In practice, the mDR-learner does not know the true nuisance parameters $g_{0}$, but estimates them, $\hat{g}$. Consequently, the upper bound easily obtained does not match our desired bound, but is instead written in terms of estimated nuisance functions:
$L(\hat{\theta},\hat{g})-L(\theta^{*},\hat{g})\leq\text{Rate}_{\theta}(\hat{g}).$
To correct for this, we turn to the criteria outlined by Foster and Syrgkanis [2023]. We begin by defining the nuisance/target algorithms and their respective rates:
Definition 1 (Algorithms & Rates).
  • a)
    Nuisance algorithms and rates. For the mDR-learner, as $g=(\pi,\mu^{0},\mu^{1},G)$, we define first stage algorithms and rates for each component of $g$.
  • The algorithm $\text{Alg}_{\pi}$ outputs a predictor of $\pi$, $\hat{\pi}$, for which
$\|\hat{\pi}-\pi_{0}\|\leq\text{Rate}_{\pi}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
  • The algorithm $\text{Alg}_{G}$ outputs a predictor of $G$, $\hat{G}$, for which
$\|\hat{G}-G_{0}\|\leq\text{Rate}_{G}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
  • $\textit{The algorithm }\text{Alg}_{\mu^{0}}$ outputs a predictor of $\mu^{0}$, $\hat{\mu}^{0}$ for which
$\|\hat{\mu}^{0}-\mu^{0}_{0}\|\leq\text{Rate}_{\mu^{0}}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
  • The algorithm $\text{Alg}_{\mu^{1}}$ outputs a predictor of $\mu^{1}$, $\hat{\mu}^{1}$ for which
$\|\hat{\mu}^{1}-\mu^{1}_{0}\|\leq\text{Rate}_{\mu^{1}}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
  • b)
    Target algorithm and rate. Let $\theta^{*}\in\hat{\Theta}$ and define the second stage algorithm $\text{Alg}(g)$, with $g=(\pi,\mu^{0},\mu^{1},G)$, to be a predictor of $\theta$, for which
$L(\hat{\theta},g)-L(\theta^{*},g)\leq\text{Rate}_{\theta}(\delta;g)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
[…]
Definition 2 (Directional Derivative). Let $\mathcal{F}$ be a vector space of functions. For a functional $F:\mathcal{F}\rightarrow\mathbbm{R}$, we define the derivative operator $\left.D_{f}F(f)[h]=\frac{d}{dt}F(f+th)\right|_{t=0}$ for a pair of functions $f,h\in\mathcal{F}$. Likewise, we define $\left.D^{k}_{f}F(f)[h_{1},...,h_{k}]=\frac{\partial^{k}}{\partial t_{1}...\partial t_{k}}F(f+t_{1}h_{1}+...+t_{k}h_{k})\right|_{t_{1}=...t_{k}=0}$. When considering a functional in two arguments, e.g. $L(\theta,g)$, we write $D_{\theta}L(\theta,g)$ and $D_{g}L(\theta,g)$ to highlight which component the argument is taken with respect to.
In order to identify an excess risk upper bound that is written in terms of known nuisance functions, Foster and Syrgkanis [2023] requires that our estimator meets four assumptions:
Assumption 1 (Orthogonal Loss). The population risk $L(\theta^{*},g_{0})$ is Neyman orthogonal:
$D_{\pi}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},\pi-\pi_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta}$ (30)
$D_{\mu^{1}}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},\mu^{1}-\mu^{1}_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta}$ (31)
$D_{\mu^{0}}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},\mu^{0}-\mu^{0}_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta}$ (32)
$D_{G}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},G-G_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta}$ (33)
Notation. For a subset $\mathcal{X}$ of a vector space, conv($\mathcal{X}$) will denote the convex hull. For an element $x\in\mathcal{X}$, we define the star hull as in (34) and adopt the shorthand star($\mathcal{X}$) :=assign:=:= star($\mathcal{X}$,0).
$\text{star}(\mathcal{X},x)=\left\{t\cdot x+(1-t)\cdot x^{\prime}|x^{\prime}\in\mathcal{X},t\in[0,1]\right\}$ (34)
Assumption 2 (First order optimality).
The minimiser for the population risk satisfies the first-order optimality condition:
$D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*}]\geq 0~{}~{}~{}\forall\theta\in\text{star}(\Theta,\theta^{*})$ (35)
Assumption 3 (Strong convexity in prediction).
The population risk $L$ is strongly convex with respect to the prediction: For all $\theta\in\Theta$ and $g\in\mathcal{G}$,
$D^{2}_{\theta}L(\bar{\theta},g)[\theta-\theta^{*},\theta-\theta^{*}]\geq\lambda\|\theta-\theta^{*}\|^{2}-\kappa\|g-g_{0}\|^{4}~{}~{}~{}\forall\bar{\theta}\in\text{star}(\hat{\Theta},\theta^{*})$ (36)
Assumption 4 (Higher-order smoothness).
There exist constants $\beta_{1},~{}\beta_{2}>0$ such that the following derivative bounds hold:
  1. a)
    Second-order smoothness with respect to target. For all $\theta\in\hat{\Theta}$ and all $\bar{\theta}\in\text{star}(\hat{\Theta},\theta^{*})$:
    $D^{2}_{\theta}L(\bar{\theta},g_{0})[\theta-\theta^{*},\theta-\theta^{*}]\leq\beta_{1}\|\theta-\theta^{*}\|^{2}$ (37)
  2. b)
    Higher-order smoothness. For all $\theta\in\text{star}(\hat{\Theta},\theta^{*}),~{}g\in\mathcal{G}$, and $\bar{g}\in\text{star}(\mathcal{G},g_{0})$:
    $|D_{g}^{2}D_{\theta}L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]|\leq\beta_{2}\|\theta-\theta^{*}\|\cdot\|g-g_{0}\|^{2}$ (38)
Proofs that demonstrate that each assumption holds for the mDR-learner can be found in Appendix F. When these assumptions hold, the excess risk bound under known nuisance parameters can written in terms of estimated nuisance parameters and their model rates.
Theorem 1, outlined below is an implementation of Theorem 1 presented by Foster and Syrgkanis [2023]. However, in this Theorem we contextualise their work to the mDR-learner, providing an insight into the rate robust properties of this estimator:
Theorem 1. Suppose that there is some $\theta^{*}\in argmin_{\theta}L(\theta,g_{0})$ such that Assumptions 1 to 4 hold (See Appendix F). Then the mDR-learner produces a parameter $\hat{\theta}$ such that with probability at least $1-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}$,
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})\leq$ $\frac{2}{1-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\text{Rate}_{\pi}^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\text{Rate}_{\pi}^{2}\cdot\text{Rate}^{2}_{\mu^{1}}+\frac{1}{\delta_{3}}\beta_{2,3}\text{Rate}^{2}_{\pi}\cdot\text{Rate}_{\mu^{0}}^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\text{Rate}^{2}_{\pi}\cdot\text{Rate}^{2}_{G}+\frac{1}{\delta_{5}}\beta_{2,5}\text{Rate}^{2}_{\mu^{1}}\cdot\text{Rate}^{2}_{G}+\frac{1}{\delta_{6}}\beta_{2,6}\text{Rate}^{4}_{G})$
where
$\beta_{2,1}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\hat{G}\hat{\pi}^{3}}\{Y-\hat{\mu}^{1}\}-\frac{(1-A)C}{\hat{G}(1-\hat{\pi})^{3}}\{Y-\hat{\mu}^{0}\}\right)^{2}\right]}$
$\beta_{2,2}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\hat{G}\hat{\pi}^{2}}\right)^{2}\right]}$
$\beta_{2,3}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\hat{G}(1-\hat{\pi})^{2}}\right)^{2}\right]}$
$\beta_{2,4}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\hat{G}^{2}\hat{\pi}^{2}}\{Y-\hat{\mu}^{1}\}-\frac{(1-A)C}{\hat{G}^{2}(1-\hat{\pi})^{2}}\{Y-\hat{\mu}^{0}\}\right)^{2}\right]}$
$\beta_{2,5}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\hat{G}^{2}\hat{\pi}}\right)^{2}\right]}$
$\beta_{2,6}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\hat{G}^{3}\hat{\pi}}\{Y-\hat{\mu}^{1}\}-\frac{(1-A)C}{\hat{G}^{3}(1-\hat{\pi})}\{Y-\hat{\mu}^{0}\}\right)^{2}\right]}$
The bound derived in Theorem 1 comprises of seven elements; six are products of nuisance estimates and their respective convergence rates, and one contains only the convergence rate of the estimates from the second stage pseudo-outcome model. With this, when the nuisance functions are known, the bound reduces to its first term, only depending on the convergence rate of the estimates from the second stage model. However, when the nuisance functions are estimated, the bound is influenced by the convergence rate of the nuisance estimates. We also observe a rate robustness property for the outcome model estimates in the mDR-learner, as when fast rates are observed for the propensity score and missingness model estimates, the outcome model estimates can converge at slower rates without increasing this upper bound. Similar observations were shown for the DR-learner [Foster and Syrgkanis, 2023], however, once again our bound shows the extension of this robustness to the missingness model.
Appendix F mDR-learner excess risk bound - Proofs
Assumption 1 - Proof
For notational ease, we write the pseudo-outcomes created by the mDR-learner, $\phi(\pi,\mu^{0},\mu^{1},G)$, as $\phi(g)$, where $g$ represents the set of nuisance functions $\{\pi,\mu^{0},\mu^{1},G\}$. We also re-write $\phi(g)$ as:
$\phi(g)=\frac{AC}{\pi G}\{Y-\mu^{1}\}-\frac{(1-A)C}{(1-\pi)G}\{Y-\mu^{0}\}+\mu^{1}-\mu^{0}$
We take the directional derivative of the MSE loss function when evaluated at its infimum with respect to the target parameter $\theta$:
$D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*}]$ $=\frac{\partial}{\partial t}\{\operatorname{\mathbb{E}}[(\phi(g_{0})-(\theta(X)\cdot t+(1-t)\cdot\theta^{*}(X)))^{2}]\}|_{t=0}$
$=-2\operatorname{\mathbb{E}}[(\phi(g_{0})-\theta^{*}(X))(\theta(X)-\theta^{*}(X))]$
Showing (30) for the mDR-learner:
$D_{\pi}D_{\theta}$ $L(\theta^{*},g_{0})[\theta-\theta^{*},\pi-\pi_{0}]$
$=-2\frac{\partial}{\partial t}\operatorname{\mathbb{E}}\left[\left(\frac{AC}{(\pi_{0}+t\cdot(\pi-\pi_{0}))G_{0}}\{Y-\mu^{1}_{0}\}-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\left.\frac{(1-A)C}{(1-\pi_{0}-t\cdot(\pi-\pi_{0}))G_{0}}\{Y-\mu^{0}_{0}\}+\mu^{1}_{0}-\mu^{0}_{0}-\theta^{*}(X)\right)(\theta(X)-\theta^{*}(X))\right]\right|_{t=0}$
$=2\operatorname{\mathbb{E}}\left[\left(\frac{AC}{G_{0}\pi_{0}^{2}}\{Y-\mu^{1}_{0}\}+\frac{(1-A)C}{G_{0}(1-\pi_{0})^{2}}\{Y-\mu^{0}_{0}\}\right)(\pi-\pi_{0})(\theta(X)-\theta^{*}(X))\right]$
$=2\operatorname{\mathbb{E}}\left[\left.\operatorname{\mathbb{E}}\left[\left(\frac{AC}{G_{0}\pi_{0}^{2}}\{Y-\mu^{1}_{0}\}+\frac{(1-A)C}{G_{0}(1-\pi_{0})^{2}}\{Y-\mu^{0}_{0}\}\right)(\pi-\pi_{0})(\theta(X)-\theta^{*}(X))\right|Z\right]\right]$
$=2\operatorname{\mathbb{E}}\left[\left.\operatorname{\mathbb{E}}\left[\frac{AC}{G_{0}\pi_{0}^{2}}\{Y-\mu^{1}_{0}\}+\frac{(1-A)C}{G_{0}(1-\pi_{0})^{2}}\{Y-\mu^{0}_{0}\}\right|Z\right](\pi-\pi_{0})(\theta(X)-\theta^{*}(X))\right]$
$=2\operatorname{\mathbb{E}}\left[\left(\frac{1}{G_{0}\pi_{0}^{2}}\{\operatorname{\mathbb{E}}[ACY|Z]-\operatorname{\mathbb{E}}[AC|Z]\mu^{1}_{0}\}+\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\left.\frac{1}{G_{0}(1-\pi_{0})^{2}}\{\operatorname{\mathbb{E}}[(1-A)CY|Z]-\operatorname{\mathbb{E}}[(1-A)C|Z]\mu^{0}_{0}\}\right)\right.(\pi-\pi_{0})(\theta(X)-\theta^{*}(X))\right]$
$=2\operatorname{\mathbb{E}}\left[\left(\frac{1}{G_{0}\pi_{0}^{2}}\{G_{0}\pi_{0}\mu_{0}^{1}-G_{0}\pi_{0}\mu^{1}_{0}\}+\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\left.\frac{1}{G_{0}(1-\pi_{0})^{2}}\{\operatorname{\mathbb{E}}[G_{0}(1-\pi_{0})\mu_{0}^{0}|Z]-G_{0}(1-\pi_{0})\mu^{0}_{0}\}\right)\right.(\pi-\pi_{0})(\theta(X)-\theta^{*}(X))\right]$
By conditional exchangeability and consistency
$=0$
Showing (31) for the mDR-learner:
$D_{\mu^{1}}D_{\theta}$ $L(\theta^{*},g_{0})[\theta-\theta^{*},\mu^{1}-\mu^{1}_{0}]$
$=-2\frac{\partial}{\partial t}\operatorname{\mathbb{E}}\left[\left(\frac{AC}{G_{0}\pi_{0}}\{Y-(\mu^{1}_{0}+t(\mu^{1}-\mu^{1}_{0}))\}-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\left.\frac{(1-A)C}{G_{0}(1-\pi_{0})}\{Y-\mu^{0}_{0}\}+\mu^{1}_{0}+t(\mu^{1}-\mu^{1}_{0})-\mu^{0}_{0}-\theta^{*}(X)\right)(\theta(X)-\theta^{*}(X))\right]\right|_{t=0}$
$=2\operatorname{\mathbb{E}}\left[\left(\frac{AC}{G_{0}\pi_{0}}(\mu^{1}-\mu^{1}_{0})-\mu^{1}+\mu^{1}_{0}\right)(\theta(X)-\theta^{*}(X))\right]$
$=2\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{E}}\left[\left.\left(\frac{AC}{G_{0}\pi_{0}}(\mu^{1}-\mu^{1}_{0})-\mu^{1}+\mu^{1}_{0}\right)(\theta(X)-\theta^{*}(X))\right|Z\right]\right]$
$=2\operatorname{\mathbb{E}}\left[\left(\frac{\operatorname{\mathbb{E}}[AC|Z]}{G_{0}\pi_{0}}(\mu^{1}-\mu^{1}_{0})-\mu^{1}+\mu^{1}_{0}\right)(\theta(X)-\theta^{*}(X))\right]$
$=2\operatorname{\mathbb{E}}\left[\left(\mu^{1}-\mu^{1}_{0}-\mu^{1}+\mu^{1}_{0}\right)(\theta(X)-\theta^{*}(X))\right]$
$=0$
(32) can be shown for the mDR-learner using the same approach as (31). And showing (33) for the mDR-learner:
$D_{G}D_{\theta}$ $L(\theta^{*},g_{0})[\theta-\theta^{*},G-G_{0}]$
$=$ $-2\frac{\partial}{\partial t}\operatorname{\mathbb{E}}\left[\left(\frac{AC}{(G_{0}+t(G-G_{0})\pi_{0}}\{Y-\mu^{1}_{0}\}-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\left.\frac{(1-A)C}{(G_{0}+t(G-G_{0})(1-\pi_{0})}\{Y-\mu^{0}_{0}\}+\mu^{1}_{0}-\mu^{0}_{0}-\theta^{*}(X)\right)(\theta(X)-\theta^{*}(X))\right]\right|_{t=0}$
$=$ $2\frac{\partial}{\partial t}\operatorname{\mathbb{E}}\left[\left(\frac{AC}{(G_{0}+t(G-G_{0})^{2}\pi_{0}}\{Y-\mu^{1}_{0}\}-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\left.\frac{(1-A)C}{(G_{0}+t(G-G_{0})^{2}(1-\pi_{0})}\{Y-\mu^{0}_{0}\}\right)(G-G_{0})(\theta(X)-\theta^{*}(X))\right]\right|_{t=0}$
$=$ $2\operatorname{\mathbb{E}}\left[\left(\frac{AC}{G_{0}^{2}\pi_{0}}\{Y-\mu^{1}_{0}\}-\frac{(1-A)C}{G_{0}^{2}(1-\pi_{0})}\{Y-\mu^{0}_{0}\}\right)(G-G_{0})(\theta(X)-\theta^{*}(X))\right]$
$=$ $2\operatorname{\mathbb{E}}\left[\left.\operatorname{\mathbb{E}}\left[\left(\frac{AC}{G_{0}^{2}\pi_{0}}\{Y-\mu^{1}_{0}\}+\frac{(1-A)C}{G_{0}^{2}(1-\pi_{0})}\{Y-\mu^{0}_{0}\}\right)(G-G_{0})(\theta(X)-\theta^{*}(X))\right|Z,A\right]\right]$
$=$ $2\operatorname{\mathbb{E}}\left[\left.\operatorname{\mathbb{E}}\left[\left(\frac{AC}{G_{0}^{2}\pi_{0}}\{Y-\mu^{1}_{0}\}+\frac{(1-A)C}{G_{0}^{2}(1-\pi_{0})}\{Y-\mu^{0}_{0}\}\right)\right|Z,A\right](G-G_{0})(\theta(X)-\theta^{*}(X))\right]$
$=$ $2\operatorname{\mathbb{E}}\left[\left(\frac{1}{G_{0}^{2}\pi_{0}}\{\operatorname{\mathbb{E}}[ACY|Z,A]-\operatorname{\mathbb{E}}[AC|Z,A]\mu^{1}_{0}\}\right.\right.+$
$~{}~{}~{}~{}~{}~{}~{}\left.\left.\frac{1}{G_{0}^{2}(1-\pi_{0})}\{\operatorname{\mathbb{E}}[(1-A)CY|Z,A]-\operatorname{\mathbb{E}}[(1-A)C|Z,A]\mu^{0}_{0}\}\right)(G-G_{0})(\theta(X)-\theta^{*}(X))\right]$
$=$ $2\operatorname{\mathbb{E}}\left[\left(\frac{1}{G_{0}^{2}\pi_{0}}\{G_{0}\mu^{1}_{0}-G\mu^{1}_{0}\}+\frac{1}{G_{0}^{2}(1-\pi_{0})}\{G_{0}\mu^{0}_{0}-G\mu^{0}_{0}\}\right)(G-G_{0})(\theta(X)-\theta^{*}(X))\right]$
$=$ 00 0
Assumption 2 proof Let us consider a $\theta_{0}\in\Theta$, then we get:
$D_{\theta}L(\theta^{*},g_{0})[\theta-\theta_{0}]$ $=-2\operatorname{\mathbb{E}}\left[(\phi(g_{0})-\theta_{0}(X))(\theta(X)-\theta_{0}(X))\right]$
$=-2\operatorname{\mathbb{E}}[\operatorname{\mathbb{E}}[(\phi(g_{0})-\theta_{0}(X))(\theta(X)-\theta_{0}(X))|X]]$
$=-2\operatorname{\mathbb{E}}[\operatorname{\mathbb{E}}[\phi(g_{0})-\theta_{0}(X)|X](\theta(X)-\theta_{0}(X))]$
$=-2\operatorname{\mathbb{E}}[(\operatorname{\mathbb{E}}[\phi(g_{0})|X]-\theta_{0}(X))(\theta(X)-\theta_{0}(X))]$
$=0$
As:
$\operatorname{\mathbb{E}}[\varphi(g_{0};\mathcal{O})|X]$ $=\operatorname{\mathbb{E}}\left[\left.\frac{AC}{G_{0}\pi_{0}}\{Y-\mu^{1}_{0}\}-\frac{(1-A)C}{G_{0}(1-\pi_{0})}\{Y-\mu^{0}_{0}\}+\mu^{1}_{0}-\mu^{0}_{0}\right|X\right]$
$=\operatorname{\mathbb{E}}\left[\left.\mu^{1}_{0}-\mu^{0}_{0}\right|X\right]$
[…]
$=\theta_{0}(X)$
Assumption 3 proof
$D^{2}_{\theta}L(\bar{\theta},g)[\theta-\theta^{*},\theta-\theta^{*}]$ $=\left.\frac{\partial^{2}}{\partial t_{1}\partial t_{2}}\operatorname{\mathbb{E}}[(\phi(g)-(\bar{\theta}+t_{1}(\theta-\theta^{*})+t_{2}(\theta-\theta^{*})))^{2}]\right|_{t_{1}=t_{2}=0}$
$=\left.\frac{\partial}{\partial t_{2}}-2\operatorname{\mathbb{E}}[(\phi(g)-(\bar{\theta}+t_{2}(\theta-\theta^{*})))(\theta-\theta^{*})]\right|_{t_{2}=0}$
$=2\operatorname{\mathbb{E}}[(\theta-\theta^{*})^{2}]$
Therefore, for some $\lambda>0$,
$2\operatorname{\mathbb{E}}[(\theta-\theta^{*})^{2}]\geq\lambda\|\theta-\theta^{*}\|^{2},$
hence Assumption 3 holds under the condition:
$\lambda\leq 2$
Assumption 4 proof
Proof (a)
$D^{2}_{\theta}L(\bar{\theta},g)[\theta-\theta^{*},\theta-\theta^{*}]$ $=\left.\frac{\partial^{2}}{\partial t_{1}\partial t_{2}}\operatorname{\mathbb{E}}[(\phi(g)-(\bar{\theta}+t_{1}(\theta-\theta^{*})+t_{2}(\theta-\theta^{*})))^{2}]\right|_{t_{1}=t_{2}=0}$
$=\left.\frac{\partial}{\partial t_{2}}-2\operatorname{\mathbb{E}}[(\phi(g)-(\bar{\theta}+t_{2}(\theta-\theta^{*})))(\theta-\theta^{*})]\right|_{t_{2}=0}$
$=2\operatorname{\mathbb{E}}[(\theta-\theta^{*})^{2}]$
Therefore, for some $\beta_{1}>0$,
$2\operatorname{\mathbb{E}}[(\theta-\theta^{*})^{2}]\leq\beta_{1}\|\theta-\theta^{*}\|^{2},$
hence Assumption 4a holds under the condition:
$\beta_{1}\geq 2$
Proof (b)
We need to find:
$D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]$
$=\operatorname{\mathbb{E}}\left[\begin{pmatrix}\pi-\pi_{0}&\mu^{1}-\mu^{1}_{0}&\mu^{0}-\mu_{0}&G-G_{0}\end{pmatrix}\nabla^{2}_{gg}\nabla_{\theta}l(\theta^{*},\bar{g})\begin{bmatrix}\pi-\pi_{0}\\ \mu^{1}-\mu^{1}_{0}\\ \mu^{0}-\mu^{0}_{0}\\ G-G_{0}\end{bmatrix}(\theta-\theta^{*})\right]$
For the mDR-learner, $\nabla_{gg}^{2}\nabla_{\theta}l(\theta^{*},\bar{g})$ takes the following form:
$\nabla_{gg}^{2}\nabla_{\theta}l(\theta^{*},\bar{g})=\begin{bmatrix}\nabla_{\pi}\nabla_{\pi}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{\pi}\nabla_{\mu^{1}}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{\pi}\nabla_{\mu^{0}}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{\pi}\nabla_{G}\nabla_{\theta}l(\theta,\bar{g})\\ \nabla_{\mu^{1}}\nabla_{\pi}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{\mu^{1}}\nabla_{\mu^{1}}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{\mu^{1}}\nabla_{\mu^{0}}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{\mu^{1}}\nabla_{G}\nabla_{\theta}l(\theta,\bar{g})\\ \nabla_{\mu^{0}}\nabla_{\pi}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{\mu^{0}}\nabla_{\mu^{1}}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{\mu^{0}}\nabla_{\mu^{0}}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{\mu^{0}}\nabla_{G}\nabla_{\theta}l(\theta,\bar{g})\\ \nabla_{G}\nabla_{\pi}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{G}\nabla_{\mu^{1}}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{G}\nabla_{\mu^{0}}\nabla_{\theta}l(\theta,\bar{g})&\nabla_{G}\nabla_{G}\nabla_{\theta}l(\theta,\bar{g})\end{bmatrix}$
When substituting in the respective derivatives we get:
$\nabla_{gg}^{2}\nabla_{\theta}l(\theta^{*},g)=$
$\begin{bmatrix}-4\left(\frac{ac}{G\pi^{3}}\{y-\mu^{1}\}-\frac{(1-a)c}{G(1-\pi)^{3}}\{y-\mu^{0}\}\right)&-\frac{2ac}{G\pi^{2}}&-\frac{2(1-a)c}{G(1-\pi)^{2}}&-2\left(\frac{ac}{G^{2}\pi^{2}}\{y-\mu^{1}\}-\frac{(1-a)c}{G^{2}(1-\pi)^{2}}\{y-\mu^{0}\}\right)\\ -\frac{2ac}{G\pi^{2}}&0&0&-\frac{2ac}{G^{2}\pi}\\ -\frac{2(1-a)c}{G(1-\pi)^{2}}&0&0&\frac{2(1-a)c}{G^{2}\pi}\\ -2\left(\frac{ac}{G^{2}\pi^{2}}\{y-\mu^{1}\}-\frac{(1-a)c}{G^{2}(1-\pi)^{2}}\{y-\mu^{0}\}\right)&-\frac{2ac}{G^{2}\pi}&\frac{2(1-a)c}{G^{2}\pi}&-4\left(\frac{ac}{G^{3}\pi}\{y-\mu^{1}\}-\frac{(1-a)c}{G^{3}(1-\pi)}\{y-\mu^{0}\}\right)\end{bmatrix}$
And therefore,
$D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]$
$=$ $\operatorname{\mathbb{E}}\left[\{-4(\pi-\pi_{0})^{2}\left(\frac{ac}{\bar{G}\bar{\pi}^{3}}\{y-\bar{\mu}^{1}\}-\frac{(1-a)c}{\bar{G}(1-\bar{\pi})^{3}}\{y-\bar{\mu}^{0}\}\right)-\frac{4ac}{\bar{G}\bar{\pi}^{2}}(\pi-\pi_{0})(\mu^{1}-\mu^{1}_{0})\right.-$
$~{}~{}~{}~{}~{}~{}\frac{4(1-a)c}{\bar{G}(1-\bar{\pi})^{2}}(\pi-\pi_{0})(\mu^{0}-\mu^{0}_{0})-4\left(\frac{ac}{\bar{G}^{2}\bar{\pi}^{2}}\{y-\bar{\mu}^{1}\}-\frac{(1-a)c}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{y-\bar{\mu}^{0}\}\right)(\pi-\pi_{0})(G-G_{0})-$
$~{}~{}~{}~{}~{}~{}\frac{4ac}{\bar{G}^{2}\bar{\pi}}(\mu^{1}-\mu^{1}_{0})(G-G_{0})+\frac{4(1-a)c}{\bar{G}^{2}(1-\bar{\pi})}(\mu^{0}-\mu^{0}_{0})(G-G_{0})-$
$~{}~{}~{}~{}~{}~{}\left.4\left(\frac{ac}{\bar{G}^{3}\bar{\pi}}\{y-\bar{\mu}^{1}\}-\frac{(1-a)c}{\bar{G}^{3}(1-\bar{\pi})}\{y-\bar{\mu}^{0}\}\right)(G-G_{0})^{2}\}(\theta-\theta^{*})\right]$
$=$ $-4E\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)(\pi-\pi_{0})^{2}(\theta-\theta^{*})\right]-$
$4\operatorname{\mathbb{E}}\left[\frac{AC}{\bar{G}\bar{\pi}^{2}}(\pi-\pi_{0})(\mu^{1}-\mu^{1}_{0})(\theta-\theta^{*})\right]-4\operatorname{\mathbb{E}}\left[\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}(\pi-\pi_{0})(\mu^{0}-\mu^{0}_{0})(\theta-\theta^{*})\right]-$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)(\pi-\pi_{0})(G-G_{0})(\theta-\theta^{*})\right]-$
$4\operatorname{\mathbb{E}}\left[\frac{AC}{\bar{G}^{2}\bar{\pi}}(\mu^{1}-\mu^{1}_{0})(G-G_{0})(\theta-\theta^{*})\right]+4\operatorname{\mathbb{E}}\left[\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})}(\mu^{0}-\mu^{0}_{0})(G-G_{0})(\theta-\theta^{*})\right]-$
$4\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)(G-G_{0})^{2}(\theta-\theta^{*})\right]$
To identify the upper bound of this, we begin by applying the Cauchy-Schwarz inequality to pull out the nuisance and target norms:
$|D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]|$
$\leq$ $4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}\|\pi-\pi_{0}\|^{2}\|\theta-\theta^{*}\|+$
$4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{2}}\right)^{2}\right]}\|\pi-\pi_{0}\|\|\mu^{1}-\mu^{1}_{0}\|\|\theta-\theta^{*}\|+$
$4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}\right)^{2}\right]}\|\pi-\pi_{0}\|\|\mu^{0}-\mu^{0}_{0}\|\|\theta-\theta^{*}\|+$
$4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}\|\pi-\pi_{0}\|\|G-G_{0}\|\|\theta-\theta^{*}\|+$
$4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}}\right)^{2}\right]}\|\mu^{1}-\mu^{1}_{0}\|\|G-G_{0}\|\|\theta-\theta^{*}\|+$
$4\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}\|G-G_{0}\|^{2}\|\theta-\theta^{*}\|$
We then re-write this as:
$|D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]|\leq 4\left(\beta_{2,1}\|\pi-\pi_{0}\|^{2}\|\theta-\theta^{*}\|+\right.$
$\beta_{2,2}\|\pi-\pi_{0}\|\|\mu^{1}-\mu^{1}_{0}\|\|\theta-\theta^{*}\|+\beta_{2,3}\|\pi-\pi_{0}\|\|\mu^{0}-\mu^{0}_{0}\|\|\theta-\theta^{*}\|+$
$\beta_{2,4}\|\pi-\pi_{0}\|\|G-G_{0}\|\|\theta-\theta^{*}\|+\beta_{2,5}\|\mu^{1}-\mu^{1}_{0}\|\|G-G_{0}\|\|\theta-\theta^{*}\|+$
$\left.\beta_{2,6}\|G-G_{0}\|^{2}\|\theta-\theta^{*}\|\right)$
Where
$\beta_{2,1}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{3}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{3}}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}$
$\beta_{2,2}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}\bar{\pi}^{2}}\right)^{2}\right]}$
$\beta_{2,3}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{(1-A)C}{\bar{G}(1-\bar{\pi})^{2}}\right)^{2}\right]}$
$\beta_{2,4}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}^{2}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{2}(1-\bar{\pi})^{2}}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}$
$\beta_{2,5}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{2}\bar{\pi}}\right)^{2}\right]}$
$\beta_{2,6}$ $=\sqrt{\operatorname{\mathbb{E}}\left[\left(\frac{AC}{\bar{G}^{3}\bar{\pi}}\{Y-\bar{\mu}^{1}\}-\frac{(1-A)C}{\bar{G}^{3}(1-\bar{\pi})}\{Y-\bar{\mu}^{0}\}\right)^{2}\right]}$
Finally, by defining $\beta_{2}=max\{\beta_{2,1},\beta_{2,2},\beta_{2,3},\beta_{2,4},\beta_{2,5},\beta_{2,6}\}$, we write the bound as:
$|D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]|\leq 4\beta_{2}\|\theta-\theta^{*}\|\cdot\|g-g_{0}\|^{2}$
Theorem 1 proof
Using second-order Taylor expansion of the risk at $\hat{g}$, there exists $\bar{\theta}\in\text{star}(\mathcal{G},\theta^{*})$ such that
$L(\hat{\theta},\hat{g})=L(\theta^{*},\hat{g})+D_{\theta}L(\theta^{*},\hat{g})[\hat{\theta}-\theta^{*}]+\frac{1}{2}D_{\theta}^{2}L(\bar{\theta},\hat{g})[\hat{\theta}-\theta^{*},\hat{\theta}-\theta^{*}]$
$\implies$ $\frac{1}{2}D_{\theta}^{2}L(\bar{\theta},\hat{g})[\hat{\theta}-\theta^{*},\hat{\theta}-\theta^{*}]=L(\hat{\theta},\hat{g})-L(\theta^{*},\hat{g})-D_{\theta}L(\theta^{*},\hat{g})[\hat{\theta}-\theta^{*}]$
Using strong convexity (Assumption 3) we get
$D_{\theta}^{2}L(\bar{\theta},\hat{g})[\hat{\theta}-\theta^{*},\hat{\theta}-\theta^{*}]\geq\lambda\|\hat{\theta}-\theta^{*}\|^{2}$
$\implies$ $\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|^{2}\leq L(\hat{\theta},\hat{g})-L(\theta^{*},\hat{g})-D_{\theta}L(\theta^{*},\hat{g})[\hat{\theta}-\theta^{*}]$
And applying the assumed rate for $\hat{\theta}$ we obtain:
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|^{2}\leq\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},\hat{g})[\hat{\theta}-\theta^{*}]$ (39)
Using Assumption 4, we are then able to apply an additional second-order Taylor expansion, which implies there exists $\bar{g}\in\text{star}(\mathcal{G},g_{0})$ such that
$-D_{\theta}L(\theta^{*},\hat{g})[\hat{\theta}-\theta^{*}]=$ $-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]-D_{g}D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*},\hat{g}-g_{0}]$
$-\frac{1}{2}D_{g}^{2}D_{\theta}L(\theta^{*},\bar{g})[\hat{\theta}-\theta^{*},\hat{g}-g_{0},\hat{g}-g_{0}]$
Which when used in combination with the orthogonality assumption (Assumption 1)
$D_{g}D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*},\hat{g}-g_{0}]=0$
We get
$-D_{\theta}L(\theta^{*},\hat{g})[\hat{\theta}-\theta^{*}]=$ $-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]-\frac{1}{2}D_{g}^{2}D_{\theta}L(\theta^{*},\bar{g})[\hat{\theta}-\theta^{*},\hat{g}-g_{0},\hat{g}-g_{0}]$
Combining this with (39) we see
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|^{2}\leq\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]-\frac{1}{2}D_{g}^{2}D_{\theta}L(\theta^{*},\bar{g})[\hat{\theta}-\theta^{*},\hat{g}-g_{0},\hat{g}-g_{0}]$
And substituting in the higher order smoothness calculation from Assumption 4b, we get
$-\frac{1}{2}D_{g}^{2}D_{\theta}$ $L(\theta^{*},\bar{g})[\hat{\theta}-\theta^{*},\hat{g}-g_{0},\hat{g}-g_{0}]\leq 2\left(\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\theta}-\theta^{*}\|+\right.$
$\beta_{2,2}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{1}-\mu^{1}_{0}\|\|\hat{\theta}-\theta^{*}\|+\beta_{2,3}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{0}-\mu^{0}_{0}\|\|\hat{\theta}-\theta^{*}\|+$
$\beta_{2,4}\|\hat{\pi}-\pi_{0}\|\|\hat{G}-G_{0}\|\|\hat{\theta}-\theta^{*}\|+\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|\|\hat{G}-G_{0}\|\|\hat{\theta}-\theta^{*}\|+$
$\left.\beta_{2,6}\|\hat{G}-G_{0}\|^{2}\|\hat{\theta}-\theta^{*}\|\right)$
And so
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|^{2}\leq$ $\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+2\left(\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\theta}-\theta^{*}\|+\right.$
$\beta_{2,2}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{1}-\mu^{1}_{0}\|\|\hat{\theta}-\theta^{*}\|+\beta_{2,3}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{0}-\mu^{0}_{0}\|\|\hat{\theta}-\theta^{*}\|+$
$\beta_{2,4}\|\hat{\pi}-\pi_{0}\|\|\hat{G}-G_{0}\|\|\hat{\theta}-\theta^{*}\|+\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|\|\hat{G}-G_{0}\|\|\hat{\theta}-\theta^{*}\|+$
$\left.\beta_{2,6}\|\hat{G}-G_{0}\|^{2}\|\hat{\theta}-\theta^{*}\|\right)$
Then using the AM-GM inequality, we observe that for any positive constants $\delta_{1}>0,\delta_{2}>0,\delta_{3}>0,\delta_{4}>0,\delta_{5}>0,\delta_{6}>0$, we have:
$\frac{\lambda}{2}\|\hat{\theta}-\theta^{*}\|^{2}\leq$ $\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|^{4}+(\delta_{1}+\delta_{2}+\delta_{3}+\delta_{4}+\delta_{5}+\delta_{6})\|\hat{\theta}-\theta^{*}\|^{2}$
Therefore, for $\delta_{1}+\delta_{2}+\delta_{3}+\delta_{4}+\delta_{5}+\delta_{6}<\frac{\lambda}{2}$ we see
$\|\hat{\theta}-\theta^{*}\|^{2}\leq$ $\frac{1}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})-D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+$
$\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{4}+\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{0}-\mu^{0}_{0}\|^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|^{4})$
And invoking Assumption 2 (35) we obtain
$\|\hat{\theta}-\theta^{*}\|^{2}\leq$ $\frac{1}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{0}-\mu^{0}_{0}\|+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|^{4})$
To identify a bound in the form of (29), we use another Taylor expansion which implies that there exists $\bar{\theta}\in\text{star}(\hat{\Theta},\theta^{*})$ such that
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})=D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+\frac{1}{2}D_{\theta}^{2}L(\bar{\theta},g_{0})[\hat{\theta}-\theta^{*},\hat{\theta}-\theta^{*}]$
The using the smoothness bound from Assumption 4a we get:
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})\leq$ $D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]+\frac{\beta_{1}}{2}\|\hat{\theta}-\theta^{*}\|^{2}$
$\leq$ $\frac{\beta_{1}}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{0}-\mu^{0}_{0}\|+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|^{4})-$
$\left(\frac{\beta_{1}}{\lambda-2(\delta_{1}+\delta_{2}+\delta_{3}+\delta_{4}+\delta_{5}+\delta_{6})}-1\right)D_{\theta}L(\theta^{*},g_{0})[\hat{\theta}-\theta^{*}]$
Then using Assumption 2 and that $\left(\frac{\beta_{1}}{\lambda-2(\delta_{1}+\delta_{2}+\delta_{3}+\delta_{4}+\delta_{5}+\delta_{6})}-1\right)\geq 0$, we get
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})\leq$ $\frac{\beta_{1}}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\|\hat{\pi}-\pi_{0}\|^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}+\frac{1}{\delta_{3}}\beta_{2,3}\|\hat{\pi}-\pi_{0}\|\|\hat{\mu}^{0}-\mu^{0}_{0}\|+$
$\frac{1}{\delta_{4}}\beta_{2,4}\|\hat{\pi}-\pi_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+\frac{1}{\delta_{5}}\beta_{2,5}\|\hat{\mu}^{1}-\mu^{1}_{0}\|^{2}\|\hat{G}-G_{0}\|^{2}+$
$\frac{1}{\delta_{6}}\beta_{2,6}\|\hat{G}-G_{0}\|^{4})$
Which can then be expressed in terms of nuisance and target rates as:
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})\leq$ $\frac{\beta_{1}}{\frac{\lambda}{2}-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\text{Rate}_{\pi}^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\text{Rate}_{\pi}^{2}\cdot\text{Rate}^{2}_{\mu^{1}}+\frac{1}{\delta_{3}}\beta_{2,3}\text{Rate}^{2}_{\pi}\cdot\text{Rate}_{\mu^{0}}^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\text{Rate}^{2}_{\pi}\cdot\text{Rate}^{2}_{G}+\frac{1}{\delta_{5}}\beta_{2,5}\text{Rate}^{2}_{\mu^{1}}\cdot\text{Rate}^{2}_{G}+\frac{1}{\delta_{6}}\beta_{2,6}\text{Rate}^{4}_{G})$
Finally, using the known conditions on $\lambda$ and $\beta_{1}$ from assumptions 3 and 4a respectively, we can minimize our bound by setting $\beta_{1}=2$ and $\lambda=2$:
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0})\leq$ $\frac{2}{1-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}}(\text{Rate}_{\theta}(\hat{g})+\frac{1}{\delta_{1}}\beta_{2,1}\text{Rate}_{\pi}^{4}+$
$\frac{1}{\delta_{2}}\beta_{2,2}\text{Rate}_{\pi}^{2}\cdot\text{Rate}^{2}_{\mu^{1}}+\frac{1}{\delta_{3}}\beta_{2,3}\text{Rate}^{2}_{\pi}\cdot\text{Rate}_{\mu^{0}}^{2}+$
$\frac{1}{\delta_{4}}\beta_{2,4}\text{Rate}^{2}_{\pi}\cdot\text{Rate}^{2}_{G}+\frac{1}{\delta_{5}}\beta_{2,5}\text{Rate}^{2}_{\mu^{1}}\cdot\text{Rate}^{2}_{G}+\frac{1}{\delta_{6}}\beta_{2,6}\text{Rate}^{4}_{G})$
Appendix G Simulation study DGP specifications
The data in the simulation study was produced using the following data structure:
$Z\sim Unif(-1,1)^{6},~{}~{}~{}A|Z\sim\text{Bernoulli}(\pi(Z)),~{}~{}~{}C|A,Z\sim\text{Bernoulli}(G(A,Z))$
$Y(0)|Z\sim\mu_{0}(Z)+\epsilon,~{}~{}~{}Y(1)|Z\sim Y(0)+\theta(Z),~{}~{}~{}\epsilon|Z\sim\mathcal{N}(0,1)$
$Y=AY(0)+(1-A)Y(1)$
Here, the treatment variable was generated using a propensity score $\pi(Z)$; the observed outcome was created using a combination of the unexposed outcome function $\mu_{0}(Z)$, the treatment effect $\theta(Z)$ and a noise parameter $\epsilon$; and the probability of this outcome being non-missing was generated using $G(A,Z)$. We also take $Z$ to be equivalent to $X$ and assume that the covariates which impact censoring occur prior to the exposure, allowing for all covariates to be adjusted for in all models.
In each DGP, 6 covariates were created, each drawn from a uniform distribution ranging from -1 to 1. The propensity score, $\pi(Z)$, used within each DGP is defined as $\pi(Z)=expit(\sum_{\{1,2,3\}}(Z_{i}^{2}-sin(3*Z_{i}))/1.5)$, and each DGP uses a combination of the following censoring, unexposed outcome and CATE functions. The complex unexposed outcome function was defined as $\mu^{0}(Z)=\sum_{\{3,4\}}(Z_{i}/2+4\mathbbm{1}\{Z_{i}>0.5\}+sin(4Z_{i}))/2$, while the simple unexposed outcome function was defined as $\mu^{0}(Z)=Z_{3}+Z_{4}$. The complex CATE was defined as $\theta(X)=\mathbbm{1}\{|Z_{3}|>0.5\}+\mathbbm{1}\{|Z_{4}|>0.5\}$, while the simple CATE was defined as $\theta(X)=0$. Then two censoring functions $G(A,Z)=expit(3-2.75A\sum_{\{4,5\}}\mathbbm{1}(|Z_{i}|>0.5))$ and $G(Z)=expit(3-2.75\sum_{\{4,5\}}\mathbbm{1}(|Z_{i}|>0.5))$ were defined, with the first imposing strong censoring only on the exposed arm, and the second imposing strong censoring evenly across both arms.
The combination of functions for each DGP can been seen in the table below:
Table 2: Complexity of DGP functions
DGP $\pi$ $G$ $\mu^{0}$ $\tau$
1 Complex Complex (Depends on A) Simple Complex
2 Complex Complex (Does not depend on A) Simple Complex
3 Complex Complex (Does not depend on A) Complex Simple
Appendix H Simulation study - Super Learner algorithm inclusion
Table 3: Super learner algorithm libraries
Algorithm Tuning parameters
Mean (SL.mean) .
Linear model (SL.lm) .
LASSO/Elastic net (SL.glmnet) nlambda = (100,250)
alpha = (0.5,1)
Random forest (SL.ranger) mtry = (1,2)
min.node.size = (10,20,50)
Neural nets (SL.nnet) size = (1,2,5)
SVM (SL.svm) nu = 1
type = C-classification
Kernel KNN (SL.kernelKnn) h = (0.05,0.25)
Appendix I Root mean square median error (RMSME)
As the DR-learner and mDR-learner use IPTWs/IPCWs, extreme propensity score/censoring probability estimates can lead to large, highly bias CATE estimates for certain individuals. Within the simulation study, this only occurs within a small subset of the simulations and consequently, the distribution of RMSE estimates ends up skewed. For this reason, a metric such as the mean RMSE can poorly represent the performance of these estimators.
Alternatively, we compare the estimators within the simulation study using the root mean square median error (RMSME). We calculate this metric by identifying the error in individual $j$’s CATE estimate for simulation $i$, $\epsilon_{ij}$. We then take the median of these errors across the simulations, $\epsilon_{j}^{med}=Median(\epsilon_{i,j})$. By doing so, extreme CATE estimates for each individual will be removed and we can then estimate the root mean squared error (RMSE) using the median errors. This metric has a similar principle to the median CATE implementation option that was outlined in Section 3.3, reducing the estimators sensitivity to extreme CATE estimates and providing a metric that ignores outlier estimates.
$RMSME=\sqrt{\frac{1}{n}\sum_{j=1}^{n}(\epsilon_{j}^{med})^{2}}$
Appendix J Additional simulation study results
Table 4: RMSME for each learner by sample size - Examples 1-3
Data Sample size mDR mEP DR EP T
Generating Available Imputation Available Imputation Available Imputation
Process Case Case Case
1 400 0.414 0.462 0.437 0.483 0.467 0.534 0.385 0.478
800 0.315 0.370 0.347 0.383 0.374 0.419 0.318 0.376
1600 0.248 0.302 0.279 0.307 0.305 0.343 0.258 0.309
3200 0.189 0.234 0.210 0.235 0.236 0.272 0.207 0.249
2 400 0.424 0.460 0.423 0.524 0.461 0.593 0.379 0.546
800 0.337 0.372 0.338 0.419 0.372 0.474 0.317 0.432
1600 0.273 0.306 0.269 0.324 0.303 0.374 0.255 0.340
3200 0.214 0.237 0.202 0.252 0.237 0.301 0.206 0.278
3 400 0.366 0.386 0.439 0.331 0.429 0.257 0.486 0.299
800 0.274 0.308 0.349 0.273 0.346 0.220 0.417 0.260
1600 0.209 0.264 0.307 0.252 0.300 0.205 0.346 0.253
3200 0.177 0.228 0.284 0.223 0.256 0.174 0.300 0.216
Figure 6: Mean RMSE for mDR-learner, mEP-learner, DR-learner, EP-learner and T-learner in three DGPs plotted by training sample size. Figures in the left column compare the mDR-learner and mEP-learner to the available case versions of the DR-learner, EP-learner and T-learner in DGP 1, 2 and 3 respectively. Figures in the right column compare the mDR-learner and mEP-learner to the DR-learner, EP-learner and T-learner when used in combination with an outcome imputation model in DGP 1, 2 and 3 respectively.
Table 5: Mean RMSE for each learner by sample size - Examples 1-3
Data Sample size mDR mEP DR EP T
Generating Available Imputation Available Imputation Available Imputation
Process Case Case Case
1 400 0.549 0.544 0.584 0.575 0.533 0.593 0.571 0.570
800 0.457 0.456 0.578 0.479 0.442 0.496 0.464 0.475
1600 0.547 0.371 0.528 0.419 0.369 0.381 0.415 0.391
3200 0.513 0.282 0.405 0.399 0.285 0.293 0.466 0.292
2 400 0.632 0.545 0.529 0.650 0.534 0.659 0.507 0.643
800 0.467 0.458 0.442 0.568 0.454 0.552 0.452 0.589
1600 0.419 0.370 0.499 0.763 0.452 0.412 0.435 0.439
3200 0.412 0.289 0.383 0.540 0.289 0.322 0.395 0.335
3 400 1.124 0.677 1.068 0.882 0.686 0.554 0.935 0.828
800 0.713 0.563 0.685 0.630 0.585 0.446 1.059 0.746
1600 0.712 0.465 0.861 0.529 0.498 0.369 0.746 0.565
3200 0.765 0.391 0.468 0.397 0.385 0.261 0.528 0.416
Appendix K mDR-learner - Panel data
K.1 Setting
Consider a setting with $L$ post baseline visits, a binary treatment $A$ recorded at baseline (visit 00) and an outcome $Y$ recorded at the final visit (visit $L$). In this setting, missingness can occur at each visit $j$ for $j\in\{1,...L\}$, with the indicator for remaining non-missing at visit $j$ defined as $C_{j}$.
Then let the covariate set $Z$ contain information which is sufficient to control for all confounders between the treatment $A$ and the outcome $Y$, and also contains sufficient information for missing outcomes to be assumed MAR. Covariate information is recorded at each visit $j$ for $j\in\{0,...,L-1\}$ and is defined as $Z_{j}$. Additionally, $C_{j}$ is identified prior to the recording of covariates $Z_{j}$ and the covariate and outcome missingness histories up to visit $j$ are defined as $\bar{\textbf{Z}}_{j}$ and $\bar{\textbf{C}}_{j}$ respectively, with $\bar{\textbf{Z}}_{j}=\{Z_{0},...,Z_{j}\}$ and $\bar{\textbf{C}}_{j}=\{C_{0},...,C_{j}\}$.
We consider balanced data, with each individual following the same visit schedule, and focus on learning the CATE, $\theta(X)=\operatorname{\mathbb{E}}[Y(1)|X]-\operatorname{\mathbb{E}}[Y(0)|X]$, where $X$ are the baseline covariates in which heterogeneity is of interest, with $X\subseteq Z$.
K.1.1 Identifiability
For the CATE to be identifiable in this setting, we require A1-A5 from Section 2.3, along with additional outcome missingness assumptions at each visit, assuring that the missing outcome data can be considered MAR even when the relationship between the missingness indicator and outcome changes over time.
We define the MAR assumption for each time $t$, $t\in\{1,...,L\}$, as conditional independence between the outcome missingness indicator at time $t$, $C_{t}$, and the observed outcome $Y$ given the treatment, $A$, the baseline covariates, $Z_{0}$, and the time varying covariates up to time $t-1$; ($Y\perp\!\!\!\perp C_{t}|A,Z_{0},\bar{\textbf{Z}}_{t-1}$).
By doing so, we can write the CATE in terms of iterative expectations as
$\operatorname{\mathbb{E}}[\mu^{1}(L)-\mu^{0}(L)|X]$
where $\mu^{a}(0)=y$ and $\mu^{a}(t)=\operatorname{\mathbb{E}}[\mu^{a}(t-1)|A=a,\bar{\textbf{C}}_{L+1-t}=\textbf{1},\bar{Z}_{L-t}]$ for $t\in\{1,...,L\}$, $a\in\{0,1\}$.
K.1.2 Algorithm
Using this statistical estimand, the panel data extension to the mDR-learner can be constructed by defining pseudo-outcomes using the EIF of the MSE for the CATE in this setting (40):
$Y_{mDR-L}=\sum_{t=1}^{L}\left(\frac{(A-\hat{\pi})\prod_{k=1}^{t}C_{k}}{\hat{\pi}(1-\hat{\pi})\hat{G}(t)}\left\{\mu^{a}(L-t)-\mu^{a}(L+1-t)\right\}\right)+\hat{\mu}^{1}(L)-\hat{\mu}^{0}(L)$ (40)
where $G(0)=1$ and $G(j)=\prod_{i=1}^{j}P(C_{i}=1|A=1,\bar{\textbf{C}}_{i-1}=\textbf{1},\bar{\textbf{Z}}_{i-1})$ for $j\in(1,...,N)$.
When $L=1$, these pseudo-outcomes reduce down to the mDR-learner pseudo-outcomes from 3.1. However, when $L>1$, the pseudo-outcomes incorporate post baseline measurements through time varying missingness/outcome predictions, with CATE estimates obtained by regressing the pseudo-outcomes against $X$. We present the algorithm for the mDR-learner panel data extension using a K-fold cross-fitting procedure below.
Algorithm 5 mDR-learner - L post basline visits
1:Split the data randomly into $K$ (e.g., 10) equal sized folds of $n$ observations from $O=(A,Z_{0},C_{1}Z_{1},...,C_{L-1}Z_{L-1},C_{L}Y)$, denoted $D_{1},\ldots,D_{K}$.
2:For $j\in 1,...,K$ and using all folds $\{D_{i},i=1,\dots,K,i\neq j\}$ except $D_{j}$, train models for
[…]
$G(t)$ $=$ $\prod_{k=1}^{t}P[C_{k}=1|A=a,\bar{\textbf{C}}_{k-1}=\textbf{1},\bar{\textbf{Z}}_{k-1}],~{}for~{}t\in(1,...,L),$ (43)
[…]
$\mu^{a}(t)$ $=$ $\operatorname{\mathbb{E}}(\mu^{a}(t-1)|A=a,\bar{\textbf{C}}_{L+1-t}=\textbf{1},\bar{\textbf{Z}}_{L-t})~{}\text{for}~{}t\in(1,...,L),a\in\{0,1\}$ (45)
(conditional outcome models in treated/untreated)
3:For all individuals in $D_{j}$ ($j\in 1,...,K$), obtain predictions $\hat{\pi}$ and $\hat{G}(t),~{}\hat{\mu}^{0}(t)$ and $\hat{\mu}^{1}(t)$, $t\in\{0,...,L\}$, based on the models fitted in the remaining folds.
4:Construct the pseudo outcomes for each individual in the data using (40).
5:Regress the pseudo outcomes $Y_{mDR-L}$ on covariates $X$, and obtain predictions of $\theta(X)$:
$\hat{\theta}_{mDR-L}(X)=\hat{\operatorname{\mathbb{E}}}[Y_{mDR-L}|X]$
K.2 Implementation notes
Obtaining outcome/missingness predictions is more complicated in the panel data extension of the mDR-learner as they must incorporate post baseline measurements. Missingness predictions can be obtained by running a pooled logistic regression model, obtaining conditional outcome missingness probabilities (conditional on being non-missing up to that time), which in turn can be used to obtain each $\hat{G}(t)$. Meanwhile, estimating the outcome predictions $\hat{\mu}^{a}(t)$ for $t\in\{1,...,L\}$, $a\in\{0,1\}$ requires an iterative process, as each $\mu^{a}(t)$ depends on $\mu^{a}(t-1)$.
When $L=1$, this translates to estimating $\mu^{a}(1)=\operatorname{\mathbb{E}}[Y|A=a,C=1,Z]$ for $a\in\{0,1\}$ (i.e., the mDR-learner in Section 3.1), and is straightforward as $Y$ is observed for all individuals that remain non-missing ($C=1$). However, for $L>1$, the outcome, $\mu^{a}(t-1)$, no longer represents an observed measure.
A naive solution would be to estimate the required outcomes $\hat{\mu}^{a}(t-1)$ using the outcome model from the previous iteration. However, by doing so, errors from the previous model will propagate through this iterative process. Additionally, as the estimates of $\mu^{a}(t-1)$ are only trained using individuals who were observed up to time $N+1-t$, this population may not be representative of the population that we need to obtain outcome estimates for (i.e. people who were observed at time $N-t$). If this is not addressed, the outcome predictions for individuals who have poor representation at that time point may end up biased.
This issue has been discussed in the context of the ATE by de Aguas et al. [2024] where they use a targeted learning based estimator to account for the population shifts at each time. However, we consider this issue in the context of the CATE. In this setting we require our outcome function estimates to be conditional on a set of individual level covariates. For this reason, we adjust for the disparity in the training/target populations by running a DR-learner at each time interval. This process generates pseudo-outcomes at each interval by taking the plug-in outcome prediction and adding a weighted error to the outcome estimates, where the weight is defined using the probability of being observed up to that time. A version of this sequential DR-learner approach is implemented on https://github.com/Matt-Pryce/mDR-learner_mEP-learner.